### v2

# Profinite rigidity witnessed by Dehn fillings of cusped hyperbolic 3-manifolds

Abstract.
Any profinite isomorphism between two cusped finite-volume hyperbolic 3-manifolds carries profinite isomorphisms between their Dehn fillings. With this observation, we prove that some cusped finite-volume hyperbolic 3-manifolds are profinitely rigid among all compact, orientable 3-manifolds, through detecting their exceptional Dehn fillings. In addition, we improved a criteria for profinite rigidity of a hyperbolic knot complement or a hyperbolic-type satellite knot complement among compact, orientable 3-manifolds, through examining its characterising slopes.
We obtain the following profinitely rigid examples: the complement of the Whitehead link, Whitehead sister link, $\frac{3}{10}$ two-bridge link; specific surgeries on one component of these links; the complement of (full) twist knots $\mathcal{K}_{n}$, Eudave-Muñoz knots $K(3,1,n,0)$, Pretzel knots $P(-3,3,2n+1)$, $5_{2}$ knot; the Berge manifold, and many more.
1. Introduction
The collection of all finite quotient groups of a finitely generated group $G$, denoted by $\mathcal{C}(G)$, reflects information of the original group $G$. When $G=\pi_{1}M^{3}$ is a 3-manifold group, $\mathcal{C}(G)$ corresponds to the lattice of finite regular coverings of $M$, i.e. the collection of all finite deck transformation groups.
It is known that among finitely generated groups, the profinite completion encodes full data of finite quotients.
Definition 1.1.
Let $G$ be a group, the profinite completion of $G$ is defined as
[…]
where $N$ ranges over all finite-index normal subgroups of $G$ (the notation $\lhd_{f}$ denotes finite-index normal subgroup).
Fact 1.2.
For two finitely generated groups $G_{1}$ and $G_{2}$, $\mathcal{C}(G_{1})=\mathcal{C}(G_{2})$ if and only if $\widehat{G_{1}}\cong\widehat{G_{2}}$.
Since 3-manifolds are largely determined by their fundamental groups, it is natural to ask whether $\mathcal{C}(\pi_{1}M)$ (or equivalently, $\widehat{\pi_{1}M}$) determines the homeomorphism type of $M$.
Definition 1.3.
Let $\mathscr{M}$ be a class of compact 3-manifolds. We say that a manifold $M\in\mathscr{M}$ is profinitely rigid in $\mathscr{M}$, if for any $N\in\mathscr{M}$, $\widehat{\pi_{1}N}\cong\widehat{\pi_{1}M}$ implies $N\cong M$.
Convention 1.4.
To avoid the trivial ambiguities, in our context, we always assume that a compact 3-manifold has no boundary spheres. Indeed, the boundary spheres can be capped off by 3-balls while preserving the fundamental group.
The profinite completion of a 3-manifold group does reflect substantial information of the 3-manifold itself. In [57], the author showed that among compact orientable 3-manifolds with empty or toral boundary, the profinite completion of fundamental group determines the homeomorphism type of a 3-manifold up to finitely many possibilities.
In fact, profinite rigidity in 3-manifolds is closely related with geometrization. In a series of works, Wilton-Zalesskii [54, 55, 56] showed that among compact, orientable, irreducible 3-manifolds with empty or incompressible toral boundary, the profinite completion determines whether the 3-manifold is geometric; it further determines the geometry type within the geometric case, and determines the JSJ-decomposition in the non-geometric case. In both cases, it is crucial to study the profinite rigidity of the geometric pieces, either with or without boundary.
The questions of profinite rigidity in seven of the eight geometries proposed by Thurston have been profoundly understood. For instance, $Sol$-manifolds are profinitely “almost” rigid [26], but not profinitely rigid by [44, 20]; in fact, the profinite rigidity in $Sol$-manifolds can be reduced to a pure number-theoretic problem. [47], together with [57, Corollary 8.3] for the bounded case, provides a complete profinite classification for orientable Seifert fibered spaces. Thus, the most challenging part in profinite rigidity of 3-manifolds is the hyperbolic case.
Liu [35] first showed that the profinite completion of the fundamental group of a finite-volume hyperbolic 3-manifold (including both closed and cusped hyperbolic manifolds) determines its homeomorphism type up to finitely many possibilities. Yet, whether profinite rigidity holds generally in the hyperbolic case is still unknown. There are a few examples of profinitely rigid closed hyperbolic 3-manifolds, for instance the Weeks manifold [8], the $\mathrm{Vol(3)}$ manifold [10], and 00-surgery of the knots ${6_{2}}$ and ${6_{3}}$ in the Rolfsen table by [15]. Some families of fibered hyperbolic 3-manifolds with cusps are also known to be profinitely rigid among all compact 3-manifolds. These examples include hyperbolic once-punctured torus bundles over $S^{1}$ by [11], including the figure-eight knot complement [9]; and hyperbolic four-punctured sphere bundles over $S^{1}$ by [17], including the “magic” manifold. The proof of profinite rigidity for these families are actually based on the congruence omnipotence property of the corresponding mapping class groups.
In this paper, we mainly focus on cusped hyperbolic 3-manifolds. The main advantage of the cusped case is that we can utilize the key technique, so called peripheral $\widehat{\mathbb{Z}}^{\times}$-regularity, introduced in [57]. Based on this technique, we prove that any two profinitely isomorphic cusped hyperbolic 3-manifolds have profinitely isomorphic Dehn fillings.
Let $M$ be a compact, orientable 3-manifold with boundary components consists of tori $\partial_{1}M,\cdots,\partial_{n}M$. For each $1\leq i\leq n$, suppose $c_{i}$ is a slope on $\partial_{i}M\cong T^{2}$, and $c_{i}$ is allowed to be an “empty” slope. We denote by $M_{(c_{i})}=M_{c_{1},\cdots,c_{n}}$ the corresponding Dehn filling along these slopes, where $\partial_{i}M$ is skipped over if $c_{i}$ is an empty slope.
Theorem A.
Suppose $M$ and $N$ are orientable cusped finite-volume hyperbolic 3-manifolds, and $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$. Then there exists a homeomorphism $\Psi:\partial M\to\partial N$ such that for any boundary slopes $(c_{i})$ on $\partial M$ (allowing empty slopes), there is an isomorphism $\widehat{\pi_{1}M_{(c_{i})}}\cong\widehat{\pi_{1}N_{(\Psi(c_{i}))}}$ which respects the peripheral structure.
Corollary 1.5.
In particular, the boundary homeomorphism $\Psi$ sends exceptional slopes to exceptional slopes.
We remark that Theorem A can also be generalized to mixed manifolds with hyperbolic pieces carrying boundary components, see Theorem 3.7.
As an application of Corollary 1.5, we found infinitely many examples of profinitely rigid cusped hyperbolic 3-manifolds, including non-fibered ones, through detecting their exceptional Dehn fillings. In fact, the following results are based on a series of profound studies on exceptional Dehn fillings, including [22, 24, 6, 32, 33], which characterise certain cusped hyperbolic 3-manifolds by their extreme patterns of exceptional Dehn fillings. This enables us to narrow down the manifolds profinitely isomorphic to the specific ones to some limited candidates that can be easily distinguished from each other.
Theorem B.
The following cusped finite-volume hyperbolic 3-manifolds are profinitely rigid among all compact, orientable 3-manifolds.
  1. (1)
    The Whitehead link complement $S^{3}\setminus\mathring{N}(\mathcal{W})$ (Figure 1(a)).
  2. (2)
    The Whitehead sister link complement $S^{3}\setminus\mathring{N}(\mathcal{WS})$ (Figure 1(b)).
  3. (3)
    The complement of the two-bridge link associated to the rational number $\frac{3}{10}$, denoted by $S^{3}\setminus\mathring{N}(\mathcal{L})$ (Figure 1(c)).
  4. (4)
    The complement of the (full) twist knots $\mathcal{K}_{n}$, $n\in\mathbb{Z}\setminus\{0,1\}$ (Figure 2).
  5. (5)
    The complement of the braid knots $\mathcal{J}_{n}$ presented by $\sigma_{2}^{-1}\sigma_{1}(\sigma_{1}\sigma_{2})^{3n}$, where $n\in\mathbb{Z}\setminus\{-1,0,1\}$ (Figure 3).
  6. (6)
    The complement of the Eudave-Muñoz knots $K(3,1,n,0)$, $n\in\mathbb{Z}\setminus\{0\}$ (Figure 4(a)); in particular, $K(3,1,1,0)$ is the Pretzel knot $P(-2,3,7)$ (Figure 4(b)).
  7. (7)
    The Berge manifold, i.e. the complement of the link $L_{A}$ in $S^{3}$ (Figure 5(a)).
  8. (8)
    $5,-1,-2,-4$, and $\frac{5}{2}$-Dehn surgery on one component of the Whitehead link shown in Figure 1(a).
  9. (9)
    $n-\frac{1}{2}$-Dehn surgery on one component of the $\frac{3}{10}$-two bridge link $\mathcal{L}$ shown in Figure 1(c), where $n\in\mathbb{Z}$.
  10. (10)
    00-surgery on the $L_{B}^{(1)}$ component of the link $L_{B}$ shown in Figure 5(b).
  11. (11)
    00-surgery on the $L_{B}^{(1)}$ component and $r$-surgery on the $L_{B}^{(2)}$ component of the link $L_{B}$ (Figure 5(b)), where $r\in\mathbb{Q}\setminus\{0,4\}$.
  12. (12)
    The manifold $M_{14}$ introduced in [24], which is the double branched cover of $S^{2}\times I$ along the tangle $\mathcal{Q}$ shown in Figure 6.
(a) The Whitehead link $\mathcal{W}$
(b) The Whitehead sister link $\mathcal{WS}$
(c) The $\frac{3}{10}$-two bridge link $\mathcal{L}$
Figure 1. Link diagrams for $\mathcal{W}$, $\mathcal{WS}$, $\mathcal{L}$
(a) The twist knot $\mathcal{K}_{n}$ ($n\geq 0$)
(b) The twist knot $\mathcal{K}_{n}$ ($n<0$)
Figure 2. Knot diagrams for $\mathcal{K}_{n}$
(a) The braid knot $\mathcal{J}_{n}$
(b) Notation for full twists
Figure 3. Knot diagram for $\mathcal{J}_{n}$
(a) The Eudave-Muñoz knot $K(3,1,n,0)$
(b) $K(3,1,1,0)$ is the $(-2,3,7)$-Pretzel knot
Figure 4. Knot diagram for $K(3,1,n,0)$
(a) The Berge manifold $S^{3}\setminus\mathring{N}(L_{A})$
(b) The link $L_{B}$
Figure 5. Link diagrams for $L_{A}$ and $L_{B}$
Figure 6. The tangle $\mathcal{Q}$ in $S^{2}\times I$, [24, Figure 22.14]
Remark 1.6.
  1. (1)
    The twist knot $\mathcal{K}_{-1}$ is in fact the figure-eight knot, and the profinite rigidity of the figure-eight knot complement was originally proven by Bridson and Reid [9].
  2. (2)
    The $L_{B}^{(2)}$ and $L_{B}^{(3)}$ components of the link $L_{B}$ can be viewed as a $4$-braid in the complement of the $L_{B}^{(1)}$ component, which is homeomorphic to $D^{2}\times S^{1}$. Thus, the 00-surgery on the $L_{B}^{(1)}$ component is in fact a four-punctured sphere bundle over $S^{1}$. The profinite rigidity of this manifold then follows from a general result recently proven by Cheetham-West [17].
  3. (3)
    The integral Dehn surgeries $\mathcal{W}(-1)$, $\mathcal{W}(-2)$, $\mathcal{W}(-4)$, $\mathcal{W}(5)$ are in fact once-punctured torus bundle over circle according to [29, Proposition 3], and their profinite rigidity follows from a general result by Bridson, Reid and Wilton [11].
However, our method of proof is different from [9, 17, 11], which may provide some new insights.
Remark 1.7.
It is conjectured that the profinite completion of fundamental group of a finite-volume hyperbolic 3-manifold determines its hyperbolic volume. The examples listed in Theorem B include some cusped hyperbolic 3-manifolds with small volumes. For instance, the Whitehead link complement and the Whitehead sister link complement are exactly the minimal volume orientable hyperbolic 3-manifolds with two cusps [1], and $\mathcal{W}(\frac{5}{2})$ is one of the second smallest volume orientable hyperbolic 3-manifolds with one cusp [21].
Knot or link complements provide easily accessible examples for 3-manifolds, and the profinite properties of knot complements have long been a popular topic. Wilkes [50] showed that the fundamental group of the complement of a graph knot is profinitely rigid among all fundamental groups of compact, orientable 3-manifolds, which implies the profinite rigidity of prime graph knot complements among all knot complements in $S^{3}$. Ueki [46] showed that that if two knot complements in $S^{3}$ are profinitely isomorphic, then they have the same Alexander polynomial. Cheetham-West [16] introduced a criteria for the profinite rigidity of a hyperbolic knot complement among all knot complements in $S^{3}$, through examining the characterising property and the profinite rigidity of its 00-surgery.
In fact, profinite properties of knot complements are more computable, providing examples that lucidly illustrates our techniques. As an application of Theorem A, we are able to detect hyperbolic knot complements and hyperbolic-type satellite knot complements through the finite quotient groups, and further to match up the Dehn surgeries of a profinitely isomorphic pair of such knots.
Definition 1.8.
A satellite knot $K\subseteq S^{3}$ is hyperbolic-type if $\partial N(K)$ belongs to a hyperbolic piece in the JSJ-decomposition of the knot complement $S^{3}\setminus\mathring{N}(K)$.
For example, any satellite knot with a hyperbolic pattern is a hyperbolic-type satellite knot.
Theorem C.
Let $M=S^{3}\setminus\mathring{N}(K)$ be the complement of a hyperbolic knot, resp. a hyperbolic-type satellite knot. Suppose $N$ is a compact, orientable 3-manifold such that $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$. Then, $N\cong S^{3}\setminus\mathring{N}(K^{\prime})$ is also the complement of a hyperbolic knot, resp. a hyperbolic-type satellite knot. In addition, there exists $\sigma\in\{1,-1\}$ such that for any $r\in\mathbb{Q}\cup\{\infty\}$, the $r$-surgery of $K$ is profinitely isomorphic to the $\sigma r$-surgery of $K^{\prime}$.
Based on Theorem C, we can improve the conclusions by Cheetham-West [16] in determining the profinite rigidity of knot complements through their characterising slopes.
Definition 1.9.
Let $K\subseteq S^{3}$ be a knot. A slope $\alpha\in\mathbb{Q}$ on $\partial N(K)$ is an unoriented characterising slope for $K$, if for any knot $J\subseteq S^{3}$, the Dehn surgery $J(\alpha)$ is homeomorphic to $K(\alpha)$ implies that $J$ is isotopic to $K$ or its mirror image.
There is a slight difference between the definition of an unoriented characterising slope and the usual definition of a characterising slope. Generally, when referring to a characterising slope, it is required that the homeomorphism $J(\alpha)\cong K(\alpha)$ is orientation-preserving, see Definition 4.7 for details. However, we remove the orientation-preserving restriction here, and add a “unoriented” prefix for distinction. Regardless of this distinction, if 00 is a characerising slope of $K$, then it is also an unoriented characterising slope of $K$, see Lemma 4.8.
Corollary D.
Let $K\subseteq S^{3}$ be either a hyperbolic knot, or a hyperbolic-type satellite knot. Suppose
  1. (1)
    there is an unoriented characterising slope $\alpha\in\mathbb{Q}$ of $K$, and
  2. (2)
    the Dehn surgery $K(\alpha)$ is profinitely rigid among all closed, orientable 3-manifolds.
Then the knot complement $S^{3}\setminus\mathring{N}(K)$ is profinitely rigid among all compact, orientable 3-manifolds.
As an application of Corollary D, we show the profinite rigidity of some knot complements in $S^{3}$, based on the examples of 00-characterising slopes listed in [4, 5].
Theorem E.
Let $K\subseteq S^{3}$ be one of the following knots:
  1. (1)
    the $5_{2}$ knot in the Rolfsen table [42] (Figure 7(a));
  2. (2)
    the $15n_{43522}$ knot in the Hoste-Thistlethwaite-Weeks table [30] (Figure 7(b));
  3. (3)
    the Pretzel knots $P(-3,3,2n+1)$, where $n\in\mathbb{Z}$ (Figure 7(c));
  4. (4)
    the satellite knots $\mathcal{W}^{+}(T_{2,3},2)$ and $\mathcal{W}^{-}(T_{2,3},2)$ (Figure 8).
Then $S^{3}\setminus\mathring{N}(K)$ is profinitely rigid among all compact, orientable 3-manifolds.
(a) The $5_{2}$ knot
(b) The $15n_{43522}$ knot
(c) The Pretzel knot $P(-3,3,2n+1)$
Figure 7. The hyperbolic knots in Theorem E
(a) $\mathcal{W}^{+}(T_{2,3},2)$
(b) $\mathcal{W}^{-}(T_{2,3},2)$
Figure 8. The satellite knots in Theorem E
Remark 1.10.
Cheetham-West [16, Theorem 1.1] showed that the complement of the $5_{2}$ knot, $15n_{43522}$ knot, and Pretzel knots $P(-3,3,2n+1)$ are profinitely rigid among all knot complements in $S^{3}$. We enlarge this scope to all compact, orientable 3-manifolds. In fact, this improvement can be deduced from Theorem C.
Remark 1.11.
$\mathcal{W}(\frac{5}{2})$, $S^{3}\setminus\mathcal{K}_{n}$ ($n\in\mathbb{Z}\setminus\{0,\pm 1\}$) in Theorem B, and $S^{3}\setminus 5_{2}$, $S^{3}\setminus 15n_{43522}$, $S^{3}\setminus P(-3,3,2n+1)$ in Theorem E provide new examples of profinitely rigid but non-fibered cusped hyperbolic 3-manifolds. This demonstrates a completely different flavor in the proof of profinite rigidity of cusped hyperbolic 3-manifolds comparing to [11, 17, 52].
In fact, Theorem A, i.e. the profinite detection of Dehn filling, shows that the profinite completion of fundamental group seems to impose strong restrictions on a cusped hyperbolic 3-manifold. This may provide more confidence in the following conjecture.
Conjecture 1.12 (Part of [39, Question 9]).
Are all cusped hyperbolic 3-manifolds profinitely rigid among compact, orientable 3-manifolds?
Structure of the paper
Section 2 includes preliminary materials for profinite groups, especially the profinite completion of 3-manifold groups.
In Section 3, we establish the main technique Theorem A, which enables us to profinitely match up the Dehn fillings between a profinitely isomorphic pair of cusped hyperbolic 3-manifolds.
In Section 4, we apply this technique to knot complements in $S^{3}$ and prove Theorem C as a combination of Theorem 4.1, Theorem 4.3 and Theorem 4.5. This important observation in determining profinite rigidity of knot complements leads us to the proof of Corollary D and Theorem E in § 4.3.
Section 5 serves as a preparation for the proof of Theorem B. We show a more precise version of Corollary 1.5, that the profinite completion distinguishes specific types of exceptional Dehn fillings, including non-aspherical fillings, toroidal fillings, and aspherical fillings containing embedded Klein bottles.
Finally, in Section 6, we prove Theorem B through detecting the extremely special patterns of exceptional slopes on the listed hyperbolic manifolds. (1), (2), (3) and (12) are proved as Theorem 6.8; (4), (5) and (6) are proved as Theorem 6.12; (7) is proved in Theorem 6.14; (8) is proved in Theorem 6.5; (10) is proved in Theorem 6.20; (9) and (11) are proved in Theorem 6.25.
Acknowledgement
The author sincerely thanks Yu Huang for highly inspiring discussions in the surrounding topics. The author would also like to thank his advisor Yi Liu for useful comments.
2. Preliminaries
2.1. Profinite groups
[…]
For any two finitely generated groups $G_{1}$ and $G_{2}$, the theorem of Nikolov-Segal [38] implies that any isomorphism between $\widehat{G_{1}}$ and $\widehat{G_{2}}$ as abstract groups is indeed an isomorphism as profinte groups, i.e. a homeomorphism with respect to the product topology. Thus, we shall not distinguish between an abstract isomorphism and a continuous one in the following context, as we are mainly concerning finitely generated groups.
2.2. Residual finiteness
A group $G$ is called residually finite if the intersection of all its finite-index subgroups is trivial. It is easy to see that the canonical homomorphism $G\to\widehat{G}$ is injective if and only if $G$ is residually finite.
Hempel’s theorem for residual finiteness [28], together with the virtual Haken theorem [2], yields the following result.
Proposition 2.1 ([28, 2] ).
Suppose $M$ is a compact 3-manifold. Then the fundamental group $\pi_{1}M$ is residually finite.
2.3. Profinite detection of hyperbolicity
In our context, a finite-volume hyperbolic 3-manifold refers to a compact, orientable 3-manifold with empty or incompressible toral boundary, whose interior admits a complete Riemannian metric of constant sectional curvature $-1$.
In a series of works, Wilton-Zalesskii showed that the profinite completion detects whether a compact, orientable 3-manifold is finite-volume hyperbolic. The closed case was proven in [54], and the cusped case was proven in [55], see also [40, Theorem 4.18 and 4.20].
Moreover, when $M$ is a cusped finite-volume hyperbolic 3-manifold, [56, Proposition 3.1] showed that the conjugacy classes of peripheral subgroups in $\widehat{\pi_{1}M}$ are exactly the conjugacy classes of the maximal closed subgroups isomorphic to ^$\widehat{\mathbb{Z}}^{2}$. Therefore, the number of cusps is exactly the number of such conjugacy classes in $\widehat{\pi_{1}M}$. We conclude these properties as the following proposition.
Proposition 2.2 ([54, 55, 56] ).
Let $M$ be an orientable finite-volume hyperbolic 3-manifold. Suppose $N$ is a compact, orientable 3-manifold so that $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$, then $N$ is also a finite-volume hyperbolic 3-manifold, and $N$ has the same number of cusps as $M$. In particular, $M$ is closed if and only if $N$ is closed.
2.4. Peripheral structure
Let $M$ be a compact, orientable 3-manifold with boundary. We use the notation $\partial_{i}M$ to denote a boundary component of $M$. Up to a choice of basepoint, the inclusion induces a homomorphism $\iota_{\ast}:\pi_{1}(\partial_{i}M)\to\pi_{1}(M)$, and $\partial_{i}M$ is incompressible if and only if $\iota_{\ast}$ is injective. The image of $\iota_{\ast}$ is a conjugacy representative of the peripheral subgroup corresponding to $\partial_{i}M$. When $\partial_{i}M$ is incompressible, by a slight abuse of notation, we simply denote $\pi_{1}\partial_{i}M$ as a peripheral subgroup in $\pi_{1}M$.
Let $G$ be a group and $H$ be a subgroup of $G$. We denote by $\overline{H}$ the closure of the image of $H$ in $\widehat{G}$ through the canonical homomorphism $G\to\widehat{G}$.
Proposition 2.3.
Suppose $M$ is a compact orientable 3-manifold, and $\partial_{i}M\cong T^{2}$ is an incompressible toral boundary component. Then there is a canonical isomorphism $\overline{\pi_{1}\partial_{i}M}\cong\widehat{\pi_{1}\partial_{i}M}\cong\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}\pi_{1}\partial_{i}M$.
Proof.
According to [27] any abelian subgroup of $\pi_{1}M$ is separable. As a result, $\pi_{1}M$ induces the full profinite topology on $\pi_{1}\partial_{i}M\cong\mathbb{Z}^{2}$. The standard theories of profinite groups [41, Lemma 3.2.6] implies that $\widehat{\pi_{1}\partial_{i}M}\to\widehat{\pi_{1}M}$ is injective. Note that the image of this homomorphism is exactly $\overline{\pi_{1}\partial_{i}M}$. Thus, $\overline{\pi_{1}\partial_{i}M}\cong\widehat{\pi_{1}\partial_{i}M}$. In addition, $\pi_{1}\partial_{i}M$ is free-abelian with rank $2$. According to [57, Lemma 2.9], there is a canonical isomorphism ^$\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}\pi_{1}\partial_{i}M\cong\widehat{\mathbb{Z}}^{2}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}\partial_{i}M}$. ∎
The peripheral structure of a compact 3-manifold is given by the conjugacy classes of peripheral subgroups in the fundamental group of the 3-manifold. We extend this concept into the profinite settings.
Let $M$ and $N$ be two compact orientable 3-manifolds. For each boundary component of $M$, up to a choice of basepoints, we fix the homomorphism induced by inclusion $\iota_{i_{\ast}}:\pi_{1}(\partial_{i}M)\to\pi_{1}(M)$. Similarly, fix $\rho_{j_{\ast}}:\pi_{1}(\partial_{j}N)\to\pi_{1}(N)$.
Definition 2.4.
Let $M$ and $N$ be two compact orientable 3-manifolds. Suppose $f:\widehat{\pi_{1}M}\to\widehat{\pi_{1}N}$ is an isomorphism. We say that $f$ respects the peripheral structure if there is a one-to-one correspondence between the boundary components of $M$ and $N$, denoted by $\partial_{i}M\longleftrightarrow\partial_{\sigma(i)}N$, such that for each $i$, $f(\overline{\iota_{i_{\ast}}(\pi_{1}(\partial_{i}M))})$ is a conjugate of $\overline{\rho_{\sigma(i)_{\ast}}(\pi_{1}(\partial_{\sigma(i)}N))}$ in $\widehat{\pi_{1}N}$.
We remark that Definition 2.4 also includes compressible boundary components. In addition, when $M$ and $N$ have empty boundary, any isomorphism $f:\widehat{\pi_{1}M}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N}$ respects the peripheral structure by our definition.
Analogous to Fact 1.2, the profinite completion of fundamental group together with the peripheral structure encodes the same data as the “peripheral pair quotients” of the fundamental group. As is pointed out in [57, Theorem 2.5], $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$ by an isomorphism respecting the peripheral structure if and only if $\mathcal{C}_{\ast}(\pi_{1}(M);\iota_{1_{\ast}}(\pi_{1}(\partial_{1}M)),\cdots,\iota_{n_{\ast}}(\pi_{1}(\partial_{n}M)))=\mathcal{C}_{\ast}(\pi_{1}(N);\rho_{1_{\ast}}(\pi_{1}(\partial_{1}N)),\cdots,\rho_{n_{\ast}}(\pi_{1}(\partial_{n}N)))$.
3. Peripheral $\widehat{\mathbb{Z}}^{\times}$-regularity and correspondence of Dehn fillings
3.1. Peripheral $\widehat{\mathbb{Z}}^{\times}$-regularity
The key concept introduced in [57] is the peripheral $\widehat{\mathbb{Z}}^{\times}$-regularity, which eventually implies that a profinite isomorphism between two finite-volume hyperbolic 3-manifolds carries profinite isomorphisms between their Dehn fillings.
Recall that when $M$ is a compact, orientable 3-manifold with incompressible toral boundary, we have established a canonical isomorphism $\overline{\pi_{1}\partial_{i}M}\cong\widehat{\pi_{1}\partial_{i}M}\cong\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}\pi_{1}\partial_{i}M$ according to Proposition 2.3, where $\overline{\pi_{1}\partial_{i}M}$ denotes the closure of the peripheral subgroup in $\widehat{\pi_{1}M}$.
Definition 3.1 (Peripheral $\widehat{\mathbb{Z}}^{\times}$-regularity).
Let $M$ and $N$ be compact, orientable 3-manifolds with incompressible toral boundary. Let $\partial_{i}M$ be a boundary component of $M$ and $\pi_{1}\partial_{i}M$ be a conjugacy representative of the corresponding peripheral subgroup. Suppose $f:\widehat{\pi_{1}M}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N}$ is an isomorphism of profinite groups. We say that $f$ is peripheral $\widehat{\mathbb{Z}}^{\times}$-regular at $\partial_{i}M$ if the following holds.
  1. (1)
    There exists a boundary component $\partial_{j}N$ of $N$, a conjugacy representative of the peripheral subgroup $\pi_{1}\partial_{j}N$, and an element $g\in\widehat{\pi_{1}N}$ such that $f(\overline{\pi_{1}\partial_{i}M})=g^{-1}\overline{\pi_{1}\partial_{j}N}g$.
  2. (2)
    Let $C_{g}$ denote the conjugation by $g$, and $C_{g}\circ f$ restricts to an isomorphism
[…]
    There exist an element $\lambda\in\widehat{\mathbb{Z}}^{\times}$ and an isomorphism $\psi:\pi_{1}\partial_{i}M\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}\partial_{j}N$ such that $C_{g}\circ f=\lambda\otimes\psi$, where $\lambda$ denotes the scalar multiplication by $\lambda$.
One essential result in [57] is that any profinite isomorphism between finite-volume hyperbolic 3-manifolds is always peripheral $\widehat{\mathbb{Z}}^{\times}$-regular.
Theorem 3.2 ([57, Theorem 7.1] ).
Let $M$ and $N$ be orientable finite-volume hyperbolic 3-manifolds with cusps, and $f:\widehat{\pi_{1}M}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N}$ be an isomorphism.
  1. (1)
    $f$ respects the peripheral structure, i.e. there is a one-to-one correspondence between the boundary components of $M$ and $N$, which we simply denote as $\partial_{i}M\leftrightarrow\partial_{i}N$, such that $f(\overline{\pi_{1}\partial_{i}M})$ is a conjugate of $\overline{\pi_{1}\partial_{i}N}$ in $\widehat{\pi_{1}N}$.
  2. (2)
    $f$ is peripheral $\widehat{\mathbb{Z}}^{\times}$-regular at every boundary component of $M$:
    $C_{g_{i}}\circ f|_{\overline{\pi_{1}\partial_{i}M}}:\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}\pi_{1}\partial_{i}M\xrightarrow{\lambda_{i}\otimes\psi_{i}}\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}\pi_{1}\partial_{i}N$
    for some $g_{i}\in\widehat{\pi_{1}N}$, $\lambda_{i}\in\widehat{\mathbb{Z}}^{\times}$ and some isomorphism $\psi_{i}:\pi_{1}\partial_{i}M\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}\partial_{i}N$.
In fact, we can prove that the coefficients $\lambda_{i}$ corresponding to different boundary components are equal up to a $\pm$-sign. The proof of this theorem was actually based on Liu’s observation of homological $\widehat{\mathbb{Z}}^{\times}$-regularity [35].
More generally, peripheral $\widehat{\mathbb{Z}}^{\times}$-regularity also appears in profinite isomorphisms between mixed 3-manifolds, based on the profinite detection of JSJ-decomposition [56].
In the following context, a mixed 3-manifold refers to a compact, orientable, irreducible 3-manifold with empty or incompressible toral boundary, which admits non-trivial JSJ-decomposition and contains at least one hyperbolic JSJ-piece.
Proposition 3.3.
Let $M$ be a mixed 3-manifold, and let $\partial_{1}^{hyp}M,\cdots,\partial_{n}^{hyp}M$ be the boundary components of $M$ belonging to a hyperbolic JSJ-piece. Suppose $N$ is a compact, orientable 3-manifold, and $f:\widehat{\pi_{1}M}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N}$ is an isomorphism.
  1. (1)
    $N$ is also a mixed 3-manifold; in particular, $N$ is irreducible and boundary-incompressible.
  2. (2)
    There are exactly $n$ boundary components of $N$ which belong to hyperbolic JSJ-pieces, and we denote these boundary components as $\partial_{1}^{hyp}N,\cdots,\partial_{n}^{hyp}N$.
  3. (3)
    Up to a reordering, $f(\overline{\pi_{1}\partial_{i}^{hyp}M})$ is a conjugate of $\overline{\pi_{1}\partial_{i}^{hyp}N}$ in $\widehat{\pi_{1}N}$.
  4. (4)
    $f$ is peripheral $\widehat{\mathbb{Z}}^{\times}$-regular at each $\partial_{i}^{hyp}M$ ($1\leq i\leq n$).
The first three items are well-known for experts, and the fourth item is a direct corollary of Theorem 3.2. For completeness, we include a proof of Proposition 3.3 in Appendix A.
3.2. Dehn filling
Let us first recall the definition of Dehn filling. Let $T^{2}$ denote a torus, the collection of slopes on $T^{2}$ is denoted by
$\text{slp}(T^{2})=\left.\left\{\alpha\in\pi_{1}(T^{2})\left|\,\begin{gathered}\alpha\text{ is a non-zero primitive element in}\\ \text{the free abelian group }\pi_{1}(T^{2})\cong\mathbb{Z}^{2}\end{gathered}\right.\right\}\right/\alpha\sim-\alpha.$
For a non-zero primitive element $\alpha\in\pi_{1}(T^{2})$, we denote by $[\alpha]$ its image in $\text{slp}(T^{2})$. Usually, we simply denote an element in $\text{slp}(T^{2})$ by a lower-case letter, for example $c\in\text{slp}(T^{2})$.
We denote $\text{slp}^{\ast}(T^{2})=\text{slp}(T^{2})\cup\left\{\varnothing\right\}$, where $\varnothing$ is just a symbol used to denote “no Dehn filling”.
Definition 3.4 (Dehn filling).
Let $M$ be a compact, orientable 3-manifold with toral boundary components $\partial_{1}M,\cdots,\partial_{n}M$, and let $c_{i}\in\text{slp}^{\ast}(\partial_{i}M)$ be a choice of slopes. For each $1\leq i\leq n$, if $c_{i}\in\text{slp}(\partial_{i}M)$, we glue a solid torus $D^{2}\times S^{1}$ to $M$ along the boundary torus $\partial_{i}M$, so that the meridian of the solid torus is attached to $c_{i}$; and if $c_{i}=\varnothing$, we do nothing and skip this boundary component. The resulting compact orientable 3-manifold is called the Dehn filling of $M$ along $c_{1},\cdots,c_{n}$, and is denoted by $M_{c_{1},\cdots,c_{n}}$, or $M_{(c_{i})}$ for short.
Lemma 3.5.
Let $M$ be a compact, orientable 3-manifold with toral boundary components $\partial_{1}M,\cdots,\partial_{n}M$, and let $c_{i}\in\text{slp}^{\ast}(\partial_{i}M)$ be a choice of slopes. Up to a choice of basepoints, which may result in certain conjugations, we fix the maps $\varphi_{i}:\pi_{1}(\partial_{i}M)\to\pi_{1}(M)$. For $1\leq i\leq n$, let $\gamma_{i}$ be the image of $c_{i}$ in $\pi_{1}(M)$ through $\varphi_{i}$ if $c_{i}\in\text{slp}(\partial_{i}M)$; and let $\gamma_{i}$ be the identity element if $c_{i}=\varnothing$.
  1. (1)
    $\pi_{1}M_{(c_{i})}\cong\pi_{1}M/\langle\langle\gamma_{1},\cdots,\gamma_{n}\rangle\rangle$, where $\langle\langle\gamma_{1},\cdots,\gamma_{n}\rangle\rangle$ denotes the normal subgroup generated by $\gamma_{1},\cdots,\gamma_{n}$.
  2. (2)
    $\widehat{\pi_{1}M_{(c_{i})}}\cong\widehat{\pi_{1}M}/\overline{\langle\langle\gamma_{1},\cdots,\gamma_{n}\rangle\rangle}$, where $\overline{\langle\langle\gamma_{1},\cdots,\gamma_{n}\rangle\rangle}$ denotes the closed normal subgroup generated by $\gamma_{1},\cdots,\gamma_{n}$.
Proof.
In fact, a Dehn filling can be viewed as attaching 2-handles to the simple closed curves $c_{i}$, and then capping off the boundary spheres by 3-handles. Therefore, (1) is directly obtained from van Kampen’s theorem.
Indeed, we have a short exact sequence:
[…]
The right-exactness of the completion functor [41, Proposition 3.2.5] then implies the following exact sequence:
${\widehat{H}}$${\widehat{\pi_{1}M}}$${\widehat{\pi_{1}M_{(c_{i})}}}$${1.}$$\scriptstyle{\widehat{\iota}}$
Note that the image of $\widehat{\iota}:\widehat{H}\to\widehat{\pi_{1}M}$ is exactly the closure of $H$ in $\widehat{\pi_{1}M}$. Therefore, this implies (2). ∎
As a corollary of Theorem 3.2, we obtain the key technique Theorem A which is rephrased as follows. This enables us to derive information of Dehn filling from the finite quotient groups of a cusped hyperbolic 3-manfold.
Theorem A.
Suppose that $M$ and $N$ are orientable finite-volume hyperbolic 3-manifolds with cusps and that $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$. Then, there is a one-to-one correspondence between the cusps of $M$ and $N$ denoted by $\partial_{i}M\leftrightarrow\partial_{i}N$ ($1\leq i\leq n$), and there exist isomorphisms $\psi_{i}:\pi_{1}\partial_{i}M\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}\partial_{i}N$ between the conjugacy representatives of the peripheral subgroups such that for any slopes $c_{i}\in\text{slp}^{\ast}(\partial_{i}M)$ ($1\leq i\leq n$),
$\widehat{\pi_{1}M_{(c_{i})}}\cong\widehat{\pi_{1}N_{(\overline{\psi}_{i}(c_{i}))}}$
by an isomorphism respecting the peripheral structure. Here, $\overline{\psi}_{i}$ denotes the bijection $\text{slp}^{\ast}(\partial_{i}M)\to\text{slp}^{\ast}(\partial_{i}N)$ induced by the group isomorphism $\psi_{i}:\pi_{1}\partial_{i}M\to\pi_{1}\partial_{i}N$, and sends $\varnothing$ to $\varnothing$.
Proof.
We choose conjugacy representatives of the peripheral subgroups $\pi_{1}\partial_{i}M\subseteq\pi_{1}M$, $\pi_{1}\partial_{j}N\subseteq\pi_{1}N$, and fix a profinite isomorphism $f:\widehat{\pi_{1}M}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N}$.
According to Theorem 3.2 (1), $f$ respects the peripheral structure, so up to a reordering of boundary components, $f(\overline{\pi_{1}\partial_{i}M})$ is a conjugate of $\overline{\pi_{1}\partial_{i}N}$. In addition, Theorem 3.2 (2) further implies that up to composing with a conjugation, there exist $\lambda_{i}\in\widehat{\mathbb{Z}}^{\times}$ and $\psi_{i}:\pi_{1}\partial_{i}M\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}\partial_{i}N$ such that $conj\circ f|_{\overline{\pi_{1}\partial_{i}M}}=\lambda_{i}\otimes\psi_{i}$.
Let $\gamma_{i}\in\pi_{1}\partial_{i}M$ denote the representative element of $c_{i}$, and $\gamma_{i}=1$ if $c_{i}=\varnothing$. Then according to Lemma 3.5, $\widehat{\pi_{1}M_{(c_{i})}}\cong{\widehat{\pi_{1}M}}/{\overline{\langle\langle\gamma_{1},\cdots,\gamma_{n}\rangle\rangle}}$ and $\widehat{\pi_{1}N_{(\overline{\psi}_{i}(c_{i}))}}\cong\widehat{\pi_{1}N}/\overline{\langle\langle\psi_{1}(\gamma_{1}),\cdots,\psi_{n}(\gamma_{n})\rangle\rangle}$.
The peripheral $\widehat{\mathbb{Z}}^{\times}$-regularity implies that $f(\gamma_{i})$ is a conjugate of $\psi_{i}(\gamma_{i})^{\lambda_{i}}$ in $\widehat{\pi_{1}N}$, no matter $c_{i}\in\text{slp}(\partial_{i}M)$ or $c_{i}=\varnothing$. Since $\lambda_{i}\in\widehat{\mathbb{Z}}^{\times}$, $\overline{\langle\langle f(\gamma_{1}),\cdots,f(\gamma_{n})\rangle\rangle}=\overline{\langle\langle\psi_{1}(\gamma_{1})^{\lambda_{1}},\cdots,\psi_{n}(\gamma_{n})^{\lambda_{n}}\rangle\rangle}=\overline{\langle\langle\psi_{1}(\gamma_{1}),\cdots,\psi_{n}(\gamma_{n})\rangle\rangle}$ yield the same closed normal subgroup in $\widehat{\pi_{1}N}$. Therefore, $f$ descends to an isomorphism
[…]
Now we prove that $f_{\star}$ respects the peripheral structure. Indeed, $\partial M_{(c_{i})}$ and $\partial N_{(\overline{\psi}(c_{i}))}$ consists of the boundary components $\partial_{j}M$ and $\partial_{j}N$, where $c_{j}=\varnothing$. The corresponding peripheral subgroups are the images of $\iota_{j_{\star}}:\pi_{1}\partial_{j}M\subseteq\pi_{1}M\to\pi_{1}M/\langle\langle\gamma_{1},\cdots,\gamma_{n}\rangle\rangle\cong\pi_{1}M_{(c_{i})}$ and $\rho_{j_{\ast}}:\pi_{1}\partial_{j}N\subseteq\pi_{1}N\to\pi_{1}N/\langle\langle\psi_{1}(\gamma_{1}),\cdots,\psi_{n}(\gamma_{n})\rangle\rangle\cong\pi_{1}N_{(\overline{\psi}(c_{i}))}$ respectively. Recall that $f$ respects the peripheral structure, i.e. $f(\overline{\pi_{1}\partial_{j}M})=g_{j}\cdot\overline{\pi_{1}\partial_{j}N}\cdot g_{j}^{-1}$ for some $g_{j}\in\widehat{\pi_{1}N}$. Desending to $f_{\star}:\widehat{\pi_{1}M_{(c_{i})}}\cong\widehat{\pi_{1}N_{(\overline{\psi}_{i}(c_{i}))}}$, we obtain that $f_{\star}(\overline{\iota_{j_{\star}}(\pi_{1}\partial_{j}M)})=\Phi(g_{j})\cdot\overline{\rho_{j_{\star}}(\pi_{1}\partial_{j}N)}\cdot\Phi(g_{j})^{-1}$, where $\Phi:\widehat{\pi_{1}N}\to\widehat{\pi_{1}N}/\overline{\langle\langle\psi_{1}(\gamma_{1}),\cdots,\psi_{n}(\gamma_{n})\rangle\rangle}\cong\widehat{\pi_{1}N_{(\overline{\psi}(c_{i}))}}$ denotes the quotient homomorphism. Thus, $f_{\star}$ also respects the peripheral structure. ∎
Remark 3.6.
Theorem A also applies to non-primitive boundary slopes which yield orbifold Dehn fillings.
By Proposition 3.3, we can derive an analogous version of Theorem A for mixed 3-manifolds.
Theorem 3.7.
Let $M$ and $N$ be mixed 3-manifolds whose hyperbolic pieces carry some boundary components. Suppose that $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$. Then, there is a one-to-one correspondence between the boundary components of $M$ and $N$ belonging hyperbolic JSJ-pieces, denoted by $\partial_{i}^{hyp}M\leftrightarrow\partial_{i}^{hyp}N$ ($1\leq i\leq n$), and there exist isomorphisms $\psi_{i}:\pi_{1}\partial_{i}^{hyp}M\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}\partial_{i}^{hyp}N$ between the conjugacy representatives of the peripheral subgroups such that for any slopes $c_{i}\in\text{slp}^{\ast}(\partial_{i}^{hyp}M)$ ($1\leq i\leq n$),
$\widehat{\pi_{1}M_{(c_{i})}}\cong\widehat{\pi_{1}N_{(\overline{\psi}_{i}(c_{i}))}}.$
Proof.
The proof of this theorem is exactly the same as the first half of the proof of Theorem A, with Theorem 3.2 replaced by Proposition 3.3. ∎
4. Knot and link complements
4.1. Detection of knot and link complements
As a special application of Theorem A, let us consider a typical example of cusped finite-volume hyperbolic 3-manifold, namely the complement of a hyperbolic link in $S^{3}$. In the following context, we also include knots when we refer to links.
Theorem 4.1.
Let $M=S^{3}\setminus\mathring{N}(L)$ be the complement of a hyperbolic link $L$ in $S^{3}$ (i.e. the link complement is hyperbolic). Suppose $N$ is a compact orientable 3-manifold such that $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$, then $N$ is homeomorphic to the complement of a hyperbolic link $L^{\prime}\subseteq S^{3}$, where $L$ and $L^{\prime}$ have the same number of components.
Proof.
Suppose $L$ has $n$ components. Since $M$ is finite-volume hyperbolic, according to Proposition 2.2, $N$ is also finite-volume hyperbolic with $n$ cusps. According to Theorem A, up to reordering the boundary components of $N$, there exist bijections $\overline{\psi}_{i}:\text{slp}(\partial_{i}M)\to\text{slp}(\partial_{i}N)$ for $1\leq i\leq n$, such that $\widehat{\pi_{1}M_{(c_{i})}}\cong\widehat{\pi_{1}N_{(\overline{\psi}_{i}(c_{i}))}}$ for any $(c_{1},\cdots,c_{n})\in\text{slp}(\partial_{1}M)\times\cdots\times\text{slp}(\partial_{n}M)$.
Let $c_{i}\in\text{slp}(\partial_{i}M)$ be the meridian of the $i$-th component of $L$, which is the trivial Dehn filling slope. Then $M_{(c_{i})}\cong S^{3}$, and $\widehat{\pi_{1}M_{(c_{i})}}\cong\widehat{\pi_{1}N_{(\overline{\psi}_{i}(c_{i}))}}$ is the trivial group. Note that $\pi_{1}N_{(\overline{\psi}_{i}(c_{i}))}$ is residually finite by Proposition 2.1, so $\pi_{1}N_{(\overline{\psi}_{i}(c_{i}))}$ is also the trivial group. Since $N_{(\overline{\psi}_{i}(c_{i}))}$ is closed, the validity of the Poincaré conjecture implies that $N_{(\overline{\psi}_{i}(c_{i}))}\cong S^{3}$.
Let $L^{\prime}$ denote the link consisting of the core curves of the Dehn-filled solid tori in $N_{(\overline{\psi}_{i}(c_{i}))}=S^{3}$. Then $N\cong S^{3}\setminus\mathring{N}(L^{\prime})$ is a link complement, where $L^{\prime}$ has $n$ components, and $L^{\prime}$ is hyperbolic since $N\cong S^{3}\setminus\mathring{N}(L^{\prime})$ is finite-volume hyperbolic. ∎
Remark 4.2.
In particular, when $L$ has exactly one component, i.e. $L$ is a knot, we derive that any compact orientable 3-manifold profinitely isomorphic to a hyperbolic knot complement is also a hyperbolic knot complement.
The same procedure also holds for a link complement which is a mixed manifold, whose boundary components belong to hyperbolic pieces. For brevity of notion, we only describe this statement for hyperbolic-type satellite knots.
Theorem 4.3.
Let $K\subseteq S^{3}$ be a hyperbolic-type satellite knot. Denote $M=S^{3}\setminus\mathring{N}(K)$, and suppose $N$ is a compact orientable 3-manifold such that $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$. Then, $N$ is also homeomorphic to a knot complement $S^{3}\setminus\mathring{N}(K^{\prime})$, where $K^{\prime}$ is also a hyperbolic-type satellite knot.
Proof.
By assumption, $M$ is a mixed 3-manifold with exactly one boundary component, which belongs to a hyperbolic piece. Since $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$, Proposition 3.3 (1) implies that $N$ is also a mixed 3-manifold. According to Theorem 3.7, there exists a boundary component $\partial_{0}N$ of $N$ belonging to a hyperbolic piece, and an isomorphism $\psi:\pi_{1}\partial M\to\pi_{1}\partial_{0}N$ such that $\widehat{\pi_{1}M_{c}}\cong\widehat{\pi_{1}N_{\overline{\psi}(c)}}$ for all $c\in\text{slp}(\partial M)$.
Let $c$ be the meridian of $K$, so that $M_{c}\cong S^{3}$. Then $\widehat{\pi_{1}M_{c}}\cong\widehat{\pi_{1}N_{\overline{\psi}(c)}}$ is the trivial group, and the residual finiteness of $\pi_{1}N_{\overline{\psi}(c)}$ (Proposition 2.1) implies that $\pi_{1}N_{\overline{\psi}(c)}$ is also the trivial group. Therefore, $N_{\overline{\psi}(c)}$ is closed, since any compact orientable 3-manifold with non-empty (non-spherical) boundary has positive first Betti number. In particular, $\partial_{0}N$ is the unique boundary component of $N$. The validity of the Poincaré conjecture then implies that $N_{\overline{\psi}(c)}\cong S^{3}$.
Let $K^{\prime}$ be the core curve of the Dehn-filled solid torus in $N_{\overline{\psi}(c)}\cong S^{3}$. Then $N\cong S^{3}\setminus\mathring{N}(K^{\prime})$. Since $N$ is a mixed manifold, $K^{\prime}$ must be a satellite knot. In addition, $\partial_{0}N=\partial(N({K^{\prime}}))$ belongs to a hyperbolic piece, so $K^{\prime}$ is a hyperbolic-type satellite knot. ∎
4.2. Identifying Dehn surgery coefficients
Indeed, for the above types of knot complements, we can obtain a more precise version of Theorem A by explicitly matching up the Dehn-surgery coefficients, as shown in Theorem 4.5 below.
Lemma 4.4.
  1. (1)
    Let $M$ be a compact, orientable 3-manifold such that $\partial M$ is a torus. Then there exists a unique slope $c\in\text{slp}(\partial M)$ so that $c$ represents the null homology class in $H_{1}(M;\mathbb{Q})$; equivalently, (any representative of) $c$ belongs to the kernel of the homomorphism $\iota_{\ast}:H_{1}(\partial M;\mathbb{Z})\to H_{1}(M;\mathbb{Z})_{\mathrm{free}}$ induced by inclusion. This boundary slope is called the “rational null-homologous boundary slope” of $M$.
  2. (2)
    Suppose $M$ and $N$ are compact, orientable 3-manifolds such that $\partial M\cong\partial N\cong T^{2}$. Let $c_{0}\in\text{slp}(\partial M)$ and $c_{0}^{\prime}\in\text{slp}(\partial N)$ be the rational null-homologous boundary slopes of $M$ and $N$ respectively. Suppose $\overline{\psi}:\text{slp}(\partial M)\to\text{slp}(\partial N)$ is a bijection such that $\widehat{\pi_{1}M_{c}}\cong\widehat{\pi_{1}N_{\overline{\psi}(c)}}$ for any $c\in\text{slp}(\partial M)$. Then $\overline{\psi}(c_{0})=c_{0}^{\prime}$.
Proof.
(1) follows from a general argument using Poincaré-Lefschetz duality and the long exact sequence of homology groups, see [34, Lemma 8.15].
For (2), $\widehat{\pi_{1}M_{c}}\cong\widehat{\pi_{1}N_{\overline{\psi}(c)}}$ implies that $H_{1}(M_{c};\mathbb{Z})\cong H_{1}(N_{\overline{\psi}(c)};\mathbb{Z})$, so $b_{1}(M_{c})=b_{1}(N_{\overline{\psi}(c)})$. Note that $b_{1}(M_{c_{0}})=b_{1}(M)$, and $b_{1}(M_{c})=b_{1}(M)-1$ if $c\neq c_{0}$. Likewise, $b_{1}(N_{c_{0}^{\prime}})=b_{1}(N)$, and $b_{1}(N_{c^{\prime}})=b_{1}(N)-1$ if $c^{\prime}\neq c_{0}^{\prime}$. Comparing the first Betti number, we obtain that $b_{1}(M)=b_{1}(N)$ and $\overline{\psi}(c_{0})=c_{0}^{\prime}$. ∎
Before completing the final part of Theorem C, let us first review the concept of Dehn surgeries along a knot in $S^{3}$.
Let $K$ be a knot in $S^{3}$, and let $X_{K}=S^{3}\setminus\mathring{N}(K)$ denote its complement. Dehn fillings of $X_{K}$ can be parametrized by ${}\mathbb{Q}\cup\{\infty\}$, once we fix an orientation of $S^{3}$. To be explicit, let $[m]$ be a slope on $\partial N(K)=\partial X_{K}$, so that the simple closed curve $[m]$ bounds a disk in $N(K)$, which is called the meridian of $K$. By Alexander duality, $\partial X_{K}$ contains a unique null-homologous slope, which we denote as $[l]$, or the preferred longitude of $K$. We further choose orientations on $m$ and $l$, so that the ordered basis $(m,l)$ yields the boundary orientation of $\partial N(K)$, where $N(K)$ inherits its orientation from $S^{3}$ as a submanifold. Then, the boundary slopes of $X_{K}$ can be expressed by $[pm+ql]$, where $(p,q)$ is a pair of coprime integers. The Dehn filling of $X_{K}$ along the slope $[pm+ql]$ is also called a $\frac{p}{q}$-Dehn surgery of $K$, where $\frac{p}{q}\in\mathbb{Q}\cup\{\infty\}$, and is denoted by $K(\frac{p}{q})$. The well-definedness follows from the fact that $[pm+ql]$ and $[-pm-ql]$ represents the same boundary slope.
Theorem 4.5.
Let $K$ and $K^{\prime}$ be hyperbolic knots or hyperbolic-type satellite knots in $S^{3}$. Suppose that $\widehat{\pi_{1}X_{K}}\cong\widehat{\pi_{1}X_{K^{\prime}}}$. Then there exists $\sigma\in\{1,-1\}$, such that for any $r\in\mathbb{Q}\cup\{\infty\}$, $\widehat{\pi_{1}K({r})}\cong\widehat{\pi_{1}K^{\prime}({\sigma r})}$.
Proof.
For brevity of notation, we denote $M=X_{K}$ and $M^{\prime}=X_{K^{\prime}}$. According to Proposition 2.2, $K$ and $K^{\prime}$ are either both hyperbolic knots, or both satellite knots. By Theorem A for hyperbolic knots and Proposition 3.3 for hyperbolic-type satellite knots, there exists an isomorphism $\psi:\pi_{1}\partial M\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}\partial M^{\prime}$ such that for any $c\in\text{slp}(\partial M)$,
(4.6) $\widehat{\pi_{1}M_{c}}\cong\widehat{\pi_{1}M^{\prime}_{\overline{\psi}(c)}}.$
Let $(m,l)$ and $(m^{\prime},l^{\prime})$ be the oriented meridian-longitude pair of $K$ and $K^{\prime}$ respectively. Note that the longitude is the rational null-homologous boundary slope of a knot complement. Thus, according to Lemma 4.4 (2), $\overline{\psi}([l])=[l^{\prime}]$, i.e. $\psi(l)=\pm l^{\prime}$.
Now we show that $\psi(m)=\pm m^{\prime}$. Indeed, $M_{[m]}\cong S^{3}$, and $\widehat{\pi_{1}M^{\prime}_{\overline{\psi}[m]}}\cong\widehat{\pi_{1}M_{[m]}}$ is the trivial group. The same argument as in the proof of Theorem 4.1, i.e. the residual finiteness of the fundamental group together with the validity of Poincaré conjecture, implies that $M^{\prime}_{\overline{\psi}[m]}$ is homeomorphic to $S^{3}$. Note that $K^{\prime}$ is a non-trivial knot; according to [23, Theorem 2], the only Dehn-surgery of $K^{\prime}$ yielding $S^{3}$ is the trivial one, i.e. the $\infty$-surgery. Therefore, $\overline{\psi}[m]$ represents the same boundary slope as $[m^{\prime}]$, i.e. $\psi(m)=\pm m^{\prime}$.
As a consequence, there exist $\sigma,\delta\in\{1,-1\}$, such that $\psi(am+bl)=\delta(\sigma am^{\prime}+bl^{\prime})$ for any $a,b\in\mathbb{Z}$.
For any coprime pair of integers $(p,q)$, $K(\frac{p}{q})=M_{[pm+ql]}$, and $K^{\prime}(\sigma\frac{p}{q})=M^{\prime}_{[\sigma pm^{\prime}+ql^{\prime}]}=M^{\prime}_{[\delta(\sigma pm^{\prime}+ql^{\prime})]}=M^{\prime}_{\overline{\psi}[pm+ql]}$. Thus, $\widehat{\pi_{1}K(\frac{p}{q})}\cong\widehat{\pi_{1}K^{\prime}(\sigma\frac{p}{q})}$ by (4.6), which finishes the proof of this proposition. ∎
4.3. Criteria for profinite rigidity of a knot complement
In Definition 1.9 we have introduced the unoriented characterising slopes on a knot $K\subseteq S^{3}$. Let us first clarify its difference with the usual definition of characterising slopes.
Definition 4.7.
Let $K$ be a knot in $S^{3}$, and $\alpha\in\mathbb{Q}$ be a slope on $\partial N(K)$.
  1. (1)
    $\alpha$ is an unoriented characterising slope of $K$ if for any knot $J\subseteq S^{3}$, $J(\alpha)\cong K(\alpha)$ implies that $J$ is isotopic to $K$ or its mirror image.
  2. (2)
    $\alpha$ is a characterising slope of $K$ if for any knot $J\subseteq S^{3}$, $J(\alpha)\cong K(\alpha)$ by an orientation-preserving homeomorphism implies that $J$ is isotopic to $K$.
Generally, for $\alpha\neq 0$, neither one of unoriented-characterising and characterising directly implies the other one. However, for the 00-slope, we have the following relation.
Lemma 4.8.
Let $K$ be a knot in $S^{3}$. If 00 is a characterising slope of $K$, then 00 is also an unoriented characterising slope of $K$.
Proof.
Suppose $J$ is a knot in $S^{3}$, and there is a homeomorphism $F:J(0)\to K(0)$.
Case 1: $F$ is orientation-preserving.
In this case, by Definition 4.7 (2), $J$ is isotopic to $K$.
Case 2: $F$ is orientation-reversing.
Let $\overline{J}$ denote the mirror image of $J$. Then $\overline{J}(0)\cong J(0)$ by an orientation-reversing homeomorphism. Thus, there exists an orientation-preserving homeomorphism $\overline{J}(0)\cong K(0)$. Then, Definition 4.7 (2) implies that $\overline{J}$ is isotopic to $K$, i.e. $J$ is the mirror image of $K$. ∎
We now prove a generalized version of Corollary D. In fact, we can generalize Definition 4.7 (1) to a set of slopes.
Definition 4.9.
Let $K\subseteq S^{3}$ be a knot. A non-empty subset $\Lambda\subseteq\mathbb{Q}$, finite or infinite, is called an unoriented characterising set of slopes of $K$ if for any knot $J\subseteq S^{3}$, $J(\alpha)\cong K(\alpha)$ for all $\alpha\in\Lambda$ implies that $J$ is isotopic to $K$ or its mirror image.
Theorem 4.10 (Generalized version of Corollary D).
Let $K\subseteq S^{3}$ be either a hyperbolic knot or a hyperbolic-type satellite knot. Suppose
  1. (1)
    there is an unoriented characterising set of slopes $\Lambda$ for $K$, and
  2. (2)
    for each $\alpha\in\Lambda$, the Dehn surgery $K(\alpha)$ is profinitely rigid among all closed, orientable 3-manifolds.
Then, the knot complement $X_{K}=S^{3}\setminus\mathring{N}(K)$ is profinitely rigid among all compact, orientable 3-manifolds.
Proof.
Let $N$ be a compact, orientable 3-manifold such that $\widehat{\pi_{1}X_{K}}\cong\widehat{\pi_{1}N}$. According to Theorem 4.1 for the hyperbolic case and Theorem 4.3 for the satellite case, $N\cong X_{J}$ is also the complement of a hyperbolic knot or a hyperbolic-type satellite knot $J$. In addition, by Theorem 4.5, there exists $\sigma\in\{\pm 1\}$ such that $\widehat{\pi_{1}K(r)}\cong\widehat{\pi_{1}J(\sigma r)}$ for all $r\in\mathbb{Q}\cup\{\infty\}$. By possibly replacing $J$ with its mirror image, we may assume $\sigma=1$.
Therefore, $\widehat{\pi_{1}K(\alpha)}\cong\widehat{\pi_{1}J(\alpha)}$ for all $\alpha\in\Lambda$. Then, condition (2) implies that $K(\alpha)\cong J(\alpha)$ for all $\alpha\in\Lambda$, and condition (1) thus implies that $J$ is either equivalent or mirror image to $K$. Hence, $N\cong X_{J}\cong X_{K}$. ∎
Note that any unoriented characterising slope of $K$ is an unoriented characterising set of slopes with one element. Therefore, Theorem 4.10 does imply Corollary D.
Now we can apply Corollary D as a technique for the proof of Theorem E.
Theorem E.
Let $K\subseteq S^{3}$ be one of the following knots:
  1. (1)
    the $5_{2}$ knot (Figure 7(a));
  2. (2)
    the $15n_{43522}$ knot (Figure 7(b));
  3. (3)
    the Pretzel knots $P(-3,3,2n+1)$, where $n\in\mathbb{Z}$ (Figure 7(c));
  4. (4)
    the satellite knots $\mathcal{W}^{+}(T_{2,3},2)$ and $\mathcal{W}^{-}(T_{2,3},2)$ Figure 8.
Then $X_{K}=S^{3}\setminus\mathring{N}(K)$ is profinitely rigid among all compact, orientable 3-manifolds.
Proof.
The knots in (1), (2) and (3) are hyperbolic knots, while the knots in (4) are hyperbolic-type satellite knots. In fact, the $5_{2}$ knot is the $k3_{2}$ hyperbolic knot as tabulated in the census [13], and the $15n_{43522}$ knot is the $k8_{128}$ hyperbolic knot in the census [14]. The hyperbolicity of $P(-3,3,2n+1)$ follows from [36]. The satellite knots $\mathcal{W}^{+}(T_{2,3},2)$ and $\mathcal{W}^{-}(T_{2,3},2)$ are Whitehead-doubles of non-trivial knots, and so they are hyperbolic-type satellite knots.
Therefore, we can adopt Corollary D to prove the profinite rigidity of these knot complements. Indeed, [5, Theorem 1.1] shows that 00 is a characterising slope for $5_{2}$, and [4, Theorem 1.1] shows that 00 is a characterising slope for $15n_{43522}$, $P(-3,3,2n+1)$ or $\mathcal{W}^{\pm}(T_{2,3},2)$. Thus, by Lemma 4.8, 00 is an unoriented characterising slope for any listed knot $K$. Hence, according to Corollary D, it suffices to show that $K(0)$ is profinitely rigid among all closed, orientable 3-manifolds.
Indeed, when $K$ is the hyperbolic knot listed in (1), (2) and (3), Cheetham-West showed in the proof of [16, Theorem 1.1] that $K(0)$ is a graph manifold with JSJ-graph a single vertex and one cycle edge. Therefore, $K(0)$ is profinitely rigid in closed, orientable 3-manifolds according to the profinite classification of closed graph manifolds by Wilkes [49, Theorem A].
We claim that when $K=\mathcal{W}^{\pm}(T_{2,3},2)$, $K(0)$ is a graph manifold with JSJ-graph shown in Figure 9(a). Let $J$ denote the companion knot of $K$, which is a right-handed trefoil. The complement $X_{J}$ is a Seifert fibered space $(D^{2};(2,1),(3,1))$. Let $M_{1}$ denote the complement of the Whitehead link, then $X_{K}=M_{1}\cup_{T^{2}}X_{J}$. Note that each boundary component of $M_{1}$ already contains a null-homologous slope, corresponding to the preferred longitude of the Whitehead link. Thus, the 00-surgery on $K$ is in fact the Dehn filling corresponding to the longitude of the pattern Whitehead link. Let $\mathcal{W}(0)$ denote 00-surgery on one component of the Whitehead link. Then, $K(0)=\mathcal{W}(0)\cup_{T^{2}}X_{J}$. In fact, according to [37], $\mathcal{W}(0)$ is a graph manifold obtained from a single Seifert fibered piece $P\times S^{1}$ by gluing together two of its boundary components, where $P=S^{2}\setminus\{3\text{ disks}\}$ denotes a pair of pants. In addition, on $\partial\mathcal{W}(0)$, the slope of a regular fiber in the Seifert fiberation is exactly the meridian of the unfilled Whitehead link component. Since $K=\mathcal{W}^{\pm}(T_{2,3},2)$ is a Whitehead double of twisting number $2$, the meridian of the pattern Whitehead link corresponds to the slope $2$ on $\partial N(J)$. On the other hand, $J=T_{2,3}$, so the slope of a regular fiber of the Seifert fibered space $X_{J}$ on $\partial N(J)$ is $6$ in the ${}\mathbb{Q}\cup\{\infty\}$-parametrization. Therefore, the gluing between $\mathcal{W}(0)$ and $X_{J}$ along $T^{2}=\partial N(J)$ is not fiber-parallel. Thus, $\partial N(J)$ is a JSJ-torus of $X_{K}$ according to [3, Proposition 1.9], and the JSJ-graph of $X_{K}$ is obtained by connecting the JSJ-graphs of $\mathcal{W}(0)$ and $X_{J}$ along a single edge corresponding to $\partial N(J)$, which is exactly shown as Figure 9(a).
$P\times S^{1}$$\left(D^{2};(2,1),(3,1)\right)$
(a) The JSJ-graph of $K(0)$
$P\times S^{1}$
(b) The JSJ-graph of $\mathcal{W}(0)$
Figure 9. The JSJ-decomposition in the proof of Theorem E
In particular, $K(0)$ is a graph manifold with a non-bipartite JSJ-graph, and it follows from [49, Theorem A] that $K(0)$ is profinitely rigid among closed orientable 3-manifolds. The proof is complete by adopting Corollary D. ∎
5. Profinite detection of exceptional Dehn fillings
In this section, we focus on cusped hyperbolic 3-manifolds. Let $M$ be a finite-volume hyperbolic 3-manifold with cusps. It is known that most Dehn fillings of $M$ are also hyperbolic [45, Theorem 5.8.2]. A Dehn filling $M_{(c_{i})}$ is called exceptional if $M_{(c_{i})}$ is not a finite-volume hyperbolic 3-manifold.
In practice, we usually consider Dehn fillings on one particular boundary component $\partial_{i_{0}}M$ of $M$, and we denote
$\mathcal{E}(M,\partial_{i_{0}}M)=\left\{c\in\text{slp}(\partial_{i_{0}}M)\mid M_{c}\text{ is an exceptional Dehn filling}\right\}$
as the collection of all exceptional slopes on $\partial_{i_{0}}M$. Indeed, [45, Theorem 5.8.2] implies that only finitely many slopes on $\partial_{i_{0}}M$ yield exceptional Dehn fillings, i.e. #$\#\mathcal{E}(M,\partial_{i_{0}}M)<+\infty$.
Note that any finite-volume hyperbolic 3-manifold is aspherical and atoroidal, and thus it does not contain an embedded Klein bottle. Therefore, we consider three particular subsets of $\mathcal{E}(M,\partial_{i_{0}}M)$:
$\begin{gathered}\mathcal{E}_{s}(M,\partial_{i_{0}}M)=\left\{c\in\text{slp}(\partial_{i_{0}}M)\mid M_{c}\text{ is not aspherical}\right\};\\ \mathcal{E}_{t}(M,\partial_{i_{0}}M)=\left\{c\in\text{slp}(\partial_{i_{0}}M)\mid M_{c}\text{ is toroidal}\right\};\\ \mathcal{E}_{k}(M,\partial_{i_{0}}M)=\left\{c\in\text{slp}(\partial_{i_{0}}M)\left|\,\begin{gathered}M_{c}\text{ is aspherical, and }M_{c}\\ \text{ contains an embedded Klein bottle}\end{gathered}\right.\right\}.\end{gathered}$
Let us first prove some general results detecting these properties. These results might be familiar to experts, but we still include detailed proofs for completeness.
5.1. Detecting prime decomposition and asphericity
Lemma 5.1.
Let $M$ and $N$ be compact, orientable 3-manifolds with empty or toral boundary. The prime decomposition of $M$ and $N$ yields the following connected sum
(5.2) $\begin{gathered}M=(D^{2}\times S^{1})^{\#r_{1}}\#(S^{2}\times S^{1})^{\#r_{2}}\#M_{1}\#\cdots\#M_{s},\\ N=(D^{2}\times S^{1})^{\#r_{1}^{\prime}}\#(S^{2}\times S^{1})^{\#r_{2}^{\prime}}\#N_{1}\#\cdots\#N_{s^{\prime}},\end{gathered}$
where the $D^{2}\times S^{1}$ factors correspond to the compressible boundary components, and the $M_{i}$ or $N_{j}$ factors are irreducible and boundary-incompressible.
Suppose $f:\widehat{\pi_{1}M}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N}$ is an isomorphism. Then $r_{1}+r_{2}=r_{1}^{\prime}+r_{2}^{\prime}$, $s=s^{\prime}$, and up to a reordering, $f$ induces an isomorphism $f_{i}:\widehat{\pi_{1}M_{i}}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N_{i}}$. In addition, if $f$ respects the peripheral structure, then each $f_{i}$ also respects the peripheral structure.
Proof.
The proof mainly follows from [51, Theorem 6.22], which further requires that $M$ and $N$ are boundary-incompressible. However, we can make some adjustments to transform the boundary-compressible case to the boundary-incompressible case.
We Dehn fill the compressible boundary components of $M$ along their null-homotopic slopes, and obtain a boundary incompressible 3-manifold
$M^{+}=(S^{2}\times S^{1})^{\#(r_{1}+r_{2})}\#M_{1}\#\cdots\#M_{s}.$
In particular, the inclusion $M\hookrightarrow M^{+}$ induces an isomorphism between the fundamental groups $\pi_{1}M\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}M^{+}$.
Similarly, let
$N^{+}=(S^{2}\times S^{1})^{\#(r_{1}^{\prime}+r_{2}^{\prime})}\#N_{1}\#\cdots\#N_{s^{\prime}}$
be the Dehn-filled manifold so that $\pi_{1}N\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi N^{+}$. Then, $f$ induces an isomorphism $f^{+}:\widehat{\pi_{1}M^{+}}\cong\widehat{\pi_{1}M}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N}\cong\widehat{\pi_{1}N^{+}}$.
[51, Theorem 6.22] shows that the profinite completion of fundamental group detects the prime decomposition of a compact, orientable, boundary-incompressible 3-manifold. Thus, $r_{1}+r_{2}=r_{1}^{\prime}+r_{2}^{\prime}$, $s=s^{\prime}$, and up to a reordering, $\widehat{\pi_{1}M_{i}}\cong\widehat{\pi_{1}N_{i}}$ for each $1\leq i\leq s$.
To be explicit, up to a choice of conjugacy representatives, $\pi_{1}M_{i}$ can be viewed as a free factor of ${\pi_{1}M^{+}}$. [41, Proposition 9.1.8] implies that $\pi_{1}M^{+}$ induces the full profinite topology on $\pi_{1}M_{i}$, so the homomorphism $\widehat{\pi_{1}M_{i}}\to\widehat{\pi_{1}M^{+}}$ is injective according to [41, Lemma 3.2.6]. Thus, we can view $\widehat{\pi_{1}M_{i}}$ and $\widehat{\pi_{1}N_{i}}$ as closed subgroups of $\widehat{\pi_{1}M^{+}}$ and $\widehat{\pi_{1}N^{+}}$ respectively. [51, Theorem 6.22] actually shows that $f^{+}$ sends each $\widehat{\pi_{1}M_{i}}$ to a conjugate of $\widehat{\pi_{1}N_{i}}$ in $\widehat{\pi_{1}N^{+}}$, and we denote $f_{i}:\widehat{\pi_{1}M_{i}}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N_{i}}$ as $f^{+}$ composing with a conjugation and restricting on $\widehat{\pi_{1}M_{i}}$.
When $f$ respects the peripheral structure, we first claim that $f^{+}$ respects the peripheral structure. Indeed, when a boundary component $\partial_{k}M$ is compressible, the image of the homomorphism $\iota_{k_{\ast}}:\pi_{1}(\partial_{k}M)\to\pi_{1}(M)$ is isomorphic to $\mathbb{Z}$. Then $\overline{\iota_{k_{\ast}}(\pi_{1}(\partial_{k}M))}$ is the image of a continuous homomorphism ^$\widehat{\mathbb{Z}}\to\widehat{\pi_{1}M}$, and is hence procyclic. On the other hand, when $\partial_{k}M$ is incompressible, Proposition 2.3 implies that $\overline{\iota_{k_{\ast}}(\pi_{1}(\partial_{k}M))}\cong\widehat{\mathbb{Z}}^{2}$, which is not procyclic. The similar conclusion also holds for $N$. Since $f$ respects the peripheral structure, there is a one-to-one correspondence between the boundary components of $M$ and $N$, denoted by $\partial_{k}M\leftrightarrow\partial_{\sigma(k)}N$ such that $f(\overline{\iota_{k_{\ast}}(\pi_{1}(\partial_{k}M))})=g_{k}\cdot\overline{\rho_{\sigma(k)_{\ast}}(\pi_{1}(\partial_{\sigma(k)}M))}\cdot g_{k}^{-1}$ for some $g_{k}\in\widehat{\pi_{1}N}$. In particular, $\overline{\iota_{k_{\ast}}(\pi_{1}(\partial_{k}M))}\cong\overline{\rho_{\sigma(k)_{\ast}}(\pi_{1}(\partial_{\sigma(k)}M))}$. Distinguished by whether this group is procyclic or not, the above reasoning implies that $\partial_{\sigma(k)}N$ is incompressible if and only if $\partial_{k}M$ is incompressible. Note that the boundary of $M^{+}$ and $N^{+}$ consist of the incompressible boundary components of $M$ and $N$ respectively. For any incompressible boundary component $\partial_{k}M$ of $M$, $\iota_{k_{\ast}}^{+}:\pi_{1}(\partial_{k}M)\xrightarrow{\iota_{k_{\ast}}}\pi_{1}M\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}M^{+}$ is exactly the homomorphism induced by inclusion; and similarly, $\rho_{\sigma(k)_{\ast}}^{+}:\pi_{1}(\partial_{\sigma(k)}N)\xrightarrow{\rho_{\sigma(k)_{\ast}}}\pi_{1}N\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}N^{+}$ is induced by inclusion. Therefore, $f^{+}(\overline{\iota_{k_{\ast}}^{+}(\pi_{1}(\partial_{k}M)})=g_{k}^{+}\cdot\overline{\rho^{+}_{\sigma(k)_{\ast}}(\pi_{1}(\partial_{\sigma(k)}M))}\cdot(g_{k}^{+})^{-1}$, where $g_{k}^{+}$ denotes the image of $g_{k}$ in $\widehat{\pi_{1}N^{+}}$. This proves that $f^{+}$ respects the peripheral structure, with $\partial_{k}M\leftrightarrow\partial_{\sigma(k)}N$ for $\partial_{k}M$ incompressible.
Finally, it suffices to show that when $f^{+}$ respects the peripheral structure, each $f_{i}$ also respects the peripheral structure. Let $\partial_{k}M_{i}$ be a boundary component of $M_{i}$. Then the assumption that $f^{+}$ respects the peripheral structure implies that $f^{+}(\overline{\pi_{1}\partial_{k}M_{i}})$ is conjugate to a peripheral subgroup $\overline{\pi_{1}\partial_{l}N_{j}}\subseteq\widehat{\pi_{1}N_{j}}$ in $\widehat{\pi_{1}N^{+}}$. On the other hand, $f^{+}(\widehat{\pi_{1}M_{i}})$ is a conjugate of $\widehat{\pi_{1}N_{i}}$. Note that $\widehat{\pi_{1}N_{1}},\cdots,\widehat{\pi_{1}N_{s^{\prime}}}$ are factors of a free profinite product $\widehat{\pi_{1}N}=\widehat{\pi_{1}N_{1}}\amalg\cdots\amalg\widehat{\pi_{1}N_{s^{\prime}}}\amalg\widehat{F_{r_{1}^{\prime}+r_{2}^{\prime}}}$, so any conjugate of $\widehat{\pi_{1}N_{i}}$ intersects trivially $\widehat{\pi_{1}N_{j}}$ unless $i=j$. Therefore, $i=j$, and $f_{i}$ sends $\overline{\pi_{1}\partial_{k}M_{i}}$ to a conjugate of the closure of a peripheral subgroup in $\widehat{\pi_{1}N_{i}}$. Similarly, $f_{i}^{-1}$ sends each peripheral subgroup of $\widehat{\pi_{1}N_{i}}$ to a conjugate of a peripheral subgroup in $\widehat{\pi_{1}M_{i}}$. Indeed, this matches up a one-to-one correspondence between the boundary components of $M_{i}$ and $N_{i}$ according to [57, Proposition 6.7], which proves that $f_{i}$ respects the peripheral structure. ∎
Corollary 5.3.
Let $M$ and $N$ be compact, orientable 3-manifolds with empty or toral boundary. Suppose $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$. Then $M$ is aspherical if and only if $N$ is aspherical, except for the case that one of $M$ and $N$ is $D^{2}\times S^{1}$ while the other one is $S^{2}\times S^{1}$.
Proof.
It is easy to verify that $M$ is non-aspherical if and only if one of the following conditions holds:
  1. (i)
    $M$ has non-trivial prime decomposition;
  2. (ii)
    $M\cong S^{2}\times S^{1}$;
  3. (iii)
    $M$ has finite fundamental group.
In the first two cases, $M$ contains an essential $S^{2}$; while in the third case, $M$ is covered by $S^{3}$. Thus, it suffices to examine these three conditions.
When $M$ has non-trivial prime decomposition, according to Lemma 5.1, $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$ implies that $N$ also has non-trivial prime decomposition, so $N$ is also non-aspherical.
When $M\cong S^{2}\times S^{1}$, $\widehat{\pi_{1}N}\cong\widehat{\pi_{1}M}\cong\widehat{\mathbb{Z}}$. The residual finiteness of $\pi_{1}N$ (cf. Proposition 2.1) implies that $\pi_{1}N\cong\mathbb{Z}$, so either $N\cong S^{2}\times S^{1}$ is non-aspherical, or $N\cong D^{2}\times S^{1}$ is the exception.
When $M$ has finite fundamental group, $\widehat{\pi_{1}N}\cong\widehat{\pi_{1}M}$ is finite, and the residual finiteness of $\pi_{1}N$ (Proposition 2.1) implies that $\pi_{1}N$ is also finite, so $N$ is also spherical. ∎
5.2. Detecting toroidality
Lemma 5.4.
Let $M$ and $N$ be compact, orientable 3-manifolds with empty or toral boundary. Suppose that $f:\widehat{\pi_{1}M}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N}$ is an isomorphism respecting the peripheral structure. Then, $M$ is toroidal if and only if $N$ is toroidal.
Proof.
We denote as in (5.2) the prime decomposition of $M$ and $N$. Then Lemma 5.1 implies that $s=s^{\prime}$, and $f$ induces isomorphisms $f_{i}:\widehat{\pi_{1}M_{i}}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N_{i}}$ respecting the peripheral structure.
Note that $M$ is toroidal if and only if one of $M_{1},\cdots,M_{s}$ is toroidal, and $N$ is toroidal if and only if one of $N_{1},\cdots,N_{s^{\prime}}$ is toroidal. Thus, it suffices to consider each irreducible and boundary-incompressible prime factor, and show that $M_{i}$ is toroidal if and only if $N_{i}$ is toroidal.
According to the hyperbolization theorem [3, Theorem 1.13], $M_{i}$ is toroidal if and only if the following three conditions hold simultaneously.
  1. (i)
    $M_{i}$ is not a finite-volume hyperbolic manifold;
  2. (ii)
    $M_{i}$ is not a closed small Seifert fibered space, i.e. $M_{i}$ does not admit a Seifert fiberation over a sphere with at most 3 singular points;
  3. (iii)
    $M_{i}$ is not homeomorphic to any of the following listed Seifert fibered spaces with boundary:
    • $S^{1}\times S^{1}\times I=(S^{1}\times I;)$;
[…]
    • $P\times S^{1}=(P;)$, where $P$ denotes a pair of pants;
    • $(S^{1}\times I;(p,q))$, $p>1$, i.e. a Seifert fibered manifold with base orbifold an annulus with one singular point;
    • $(D^{2};(p_{1},q_{1}),(p_{2},q_{2}))$, $p_{1},p_{2}>1$ and $(p_{1},p_{2})\neq(2,2)$, i.e. a Seifert fibered manifold with base orbifold a disk with two singular points and having negative Euler characteristic.
It suffices to show that these three conditions hold for $M_{i}$ if and only if they hold for $N_{i}$.
The detection of condition (i) follows directly from Proposition 2.2.
Moving on to condition (ii), when $M_{i}$ is a closed small Seifert fibered space, $N_{i}$ is also closed, and it follows from [54, Theorem 8.4] that $N_{i}$ is also a Seifert fibered space. When $\pi_{1}M_{i}$ is finite, the residual finiteness of $\pi_{1}N_{i}$ (cf. Proposition 2.1) implies that $\pi_{1}N_{i}$ is also finite, and $N_{i}$ is hence a small Seifert fibered space admitting $\mathbb{S}^{3}$ geometry. When $\pi_{1}M_{i}$ is infinite, [47, Theorem 6.1] implies that either $M_{i}$ and $N_{i}$ are homeomorphic, or $M_{i}$ and $N_{i}$ both admit $\mathbb{H}^{2}\times\mathbb{R}$ geometry and share the same base orbifold. In either case, $N_{i}$ is also a closed small Seifert fibered space.
For (iii), when $M_{i}$ is a Seifert fibered space with incompressible boundary, $M_{i}$ is irreducible and Lemma 5.1 implies that $N_{i}$ is also irreducible and boundary-incompressible. It then follows from [57, Theorem 5.6] that $N_{i}$ is also Seifert fibered. Therefore, $f_{i}:\widehat{\pi_{1}M_{i}}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N_{i}}$ is a profinite isomorphism between Seifert fibered spaces with boundary, which respects the peripheral structure. [47, Theorem 6.7] then implies that $M_{i}$ and $N_{i}$ have the same base orbifolds. The list of Seifert fibered spaces in (iii) are only characterized by their base orbifolds. Therefore, if $M_{i}$ belongs to this list, then so does $N_{i}$, which finishes the proof. ∎
Remark 5.5.
It is necessary to assume that $f$ respects the peripheral structure. In fact, let $\Sigma_{0,3}$ denote a three-punctured sphere and $\Sigma_{1,1}$ denote a once-punctured torus. Then $\pi_{1}(\Sigma_{0,3}\times S^{1})\cong F_{2}\times\mathbb{Z}\cong\pi_{1}(\Sigma_{1,1}\times S^{1})$. However, $\Sigma_{0,3}\times S^{1}$ is atoroidal while $\Sigma_{1,1}\times S^{1}$ is toroidal. Indeed, these two manifolds have different numbers of boundary components, so any isomorphism $f:\widehat{\pi_{1}(\Sigma_{0,3}\times S^{1})}\to\widehat{\pi_{1}(\Sigma_{1,1}\times S^{1})}$ does not respect the peripheral structure.
5.3. Detection of embedded Klein bottles
[…]
Lemma 5.6.
Let $M$ be a compact, orientable, aspherical 3-manifold with empty or toral boundary. Then $M$ contains an embedded Klein bottle if and only if one of the following conditions hold:
  1. (i)
    $M$ is closed, admitting $Sol$-geometry, but does not fiber over $S^{1}$.
  2. (ii)
    $M$ is not a closed $Sol$-manifold, and the JSJ-decomposition [^1] We allow trivial JSJ-decomposition. of $M$ contains a Seifert fibered piece whose (one possible) base orbifold is
    • either non-orientable,
    • or contains two singular points of index $2$.
Proof.
The “if” direction is clear. Indeed, if $M$ satisfies condition (i), then $M$ is a torus semi-bundle, so $M$ contains an embedded submanifold homeomorphic to $\mathcal{N}$, and thus contains an embedded Klein bottle. If $M$ satisfies condition (ii), then $M$ contains a Seifert fibered submanifold, whose base orbifold is either a Möbius band, or a disk with two singular points of index $2$. Then this Seifert fibered submanifold is homeomorphic to $\mathcal{N}$, and thus $M$ contains an embedded Klein bottle.
Now we prove the “only if” direction. When $M$ has compressible boundary, since $M$ is aspherical, $M\cong D^{2}\times S^{1}$ and does not contain an embedded Klein bottle. Thus, we may assume that $M$ is boundary-incompressible. Suppose that $M$ contains an embedded Klein bottle $K^{2}$. Then the regular neighbourhood of $K^{2}$ in $M$ is homeomorphic to $\mathcal{N}$.
Let $T^{2}_{\mathcal{N}}$ denote the boundary torus of $\mathcal{N}$. We first claim that $T^{2}_{\mathcal{N}}$ is incompressible in $M$. Suppose $T^{2}_{\mathcal{N}}$ is compressible; since $M$ is aspherical, $T^{2}_{\mathcal{N}}$ bounds a solid torus in $M$ according to the loop theorem. Then $M=\mathcal{N}\cup_{T^{2}_{\mathcal{N}}}D^{2}\times S^{1}$. Note that $\mathcal{N}$ admits a two-fold cover $\widetilde{\mathcal{N}}\cong S^{1}\times S^{1}\times I$, where the covering map restricts to a homeomorphism from each boundary component of $\widetilde{\mathcal{N}}$ to the boundary of $\mathcal{N}$. Therefore, $M$ admits a two-fold cover $\widetilde{M}\cong D^{2}\times S^{1}\cup_{T^{2}}D^{2}\times S^{1}$, which is either $S^{2}\times S^{1}$ or a lens space. In each case, $\widetilde{M}$ is non-aspherical, so $M$ is non-aspherical, which contradicts with our assumption.
Let $M^{\prime}=M\setminus\mathring{\mathcal{N}}$. Then $\partial M^{\prime}=\partial M\sqcup T^{2}_{\mathcal{N}}$, which is incompressible.
Let us first consider two special cases. If $M^{\prime}\cong S^{1}\times S^{1}\times I$, then $M\cong\mathcal{N}$ and satisfies condition (ii). If $M^{\prime}\cong\mathcal{N}$, then $M$ is a torus semi-bundle. Thus, $M$ is either a $Sol$-manifold satisfying condition (i), or a Seifert fibered space whose base orbifold belongs to the following three possibilities:
  • $S^{2}$ with four singular points of index $2$;
  • $\mathbb{RP}^{2}$ with two singular points of index $2$;
  • a Klein bottle $K^{2}$.
Any one of these possibilities satisfies condition (ii).
Now we assume that $M^{\prime}$ is neither homeomorphic to $S^{1}\times S^{1}\times I$, nor to $\mathcal{N}$. Consider the JSJ-decomposition of $M^{\prime}$, whose collection of JSJ-tori is denoted by $\mathcal{T}^{\prime}$. Then there is a unique JSJ-piece in $M^{\prime}$ which contains the boundary component $T^{2}_{\mathcal{N}}$, and we denote this JSJ-piece as $M_{0}$.
When $M_{0}$ is a hyperbolic piece, let $\mathcal{T}=\mathcal{T}^{\prime}\cup\{T^{2}_{\mathcal{N}}\}$. Then [3, Proposition 1.9] implies that $\mathcal{T}$ is exactly the collection of JSJ-tori in $M$. Therefore, the JSJ-decomposition of $M$ contains $\mathcal{N}$ as a Seifert fibered piece, so it satisfies condition (ii).
When $M_{0}$ is a Seifert fibered piece, note that $M^{\prime}$ is not $S^{1}\times S^{1}\times I$ or $\mathcal{N}$, so the base orbifold of $M_{0}$ has negative Euler characteristic, and $M_{0}$ admits a unique Seifert fiberation. On the other hand, $\mathcal{N}$ admits two Seifert fiberations, one of which has base orbifold a Möbius band, and the other has base orbifold a disk with two singular points of index $2$.
If the regular fiber of $M_{0}$ is not parallel to any of the two regular fibers of $\mathcal{N}$ on the common boundary $T^{2}_{\mathcal{N}}$, then by [3, Proposition 1.9], $\mathcal{T}=\mathcal{T}^{\prime}\cup\{T^{2}_{\mathcal{N}}\}$ is still the JSJ-tori of $M$. Similarly, $M$ contains $\mathcal{N}$ as a JSJ-piece, and satisfies condition (ii).
If the regular fiber of $M_{0}$ is parallel to one of the two regular fibers of $\mathcal{N}$. Then, $\mathcal{T}^{\prime}$ is in fact the JSJ-tori of $M$, and $\mathcal{N}\cup_{T^{2}_{\mathcal{N}}}M_{0}$ yields an entire Seifert fibered piece in $M$. The base orbifold of $\mathcal{N}\cup M_{0}$ is obtained by gluing the base orbifold of $M_{0}$ with the base orbifold of $\mathcal{N}$ along an $S^{1}$-boundary, and the choice of the base orbifold of $\mathcal{N}$ depends on which regular fiber of $\mathcal{N}$ is parallel to that of $M_{0}$ on $T^{2}_{\mathcal{N}}$. Thus, the base orbifold of $\mathcal{N}\cup M_{0}$ either contains a Möbius band, which is non-orientable, or contains a disk with two singular points of index $2$. This shows that $M$ satisfies condition (ii). ∎
Lemma 5.7.
Let $M$ and $N$ be orientable, aspherical Seifert fibered spaces with either empty or incompressible boundary, such that $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$. Suppose $M$ admits a Seifert fiberation over a base orbifold $\mathcal{O}$, where $\mathcal{O}$ is either non-orientable, or contains two singular points with index $2$. Then $N$ also admits a Seifert fiberation over a base orbifold $\mathcal{O}^{\prime}$, which is either non-orientable, or contains two singular points with index $2$.
Proof.
We first consider the case when $M$ is closed. Then [40, Corollary 4.12] implies that $N$ is also closed. According to [47, Theorem 6.1], if $M$ admits $\mathbb{E}^{3}$, $Nil$ or $\widetilde{\mathrm{SL}(2,\mathbb{R})}$ geometry, then $N$ is homeomorphic to $M$, and the conclusion is trivial. On the other hand, if $M$ admits $\mathbb{H}^{2}\times\mathbb{R}$ geometry, then $M$ admits a unique Seifert fiberation. [47, Theorem 6.2] implies that $N$ admits the same geometry, and has the same base orbifold as $M$, even though $M$ and $N$ may not be homeomorphic, which implies our conclusion.
Now let us consider the case when $M$ has non-empty boundary. Then the above reasoning shows that $N$ also has non-empty boundary. If $M$ is homeomorphic to $S^{1}\times S^{1}\times I$ or $\mathcal{N}$, then it is easy to verify that $N\cong M$ (cf. [57, Proposition 8.2]), and the conclusion follows directly. Thus, we may assume that $M$ fibers over a hyperbolic orbifold. In this case, $M$ admits a unique Seifert fiberation, and we denote the base orbifold as $\mathcal{O}_{M}$. In addition, $N$ also fibers over a hyperbolic orbifold, and we denote the unique base orbifold as $\mathcal{O}_{N}$. In this case, the base orbifolds $\mathcal{O}_{M}$ and $\mathcal{O}_{N}$ may not be identical. However, in [57, Proposition 8.2], it is proven that the isomorphism type of $\widehat{\pi_{1}M}$ (resp. $\widehat{\pi_{1}N}$) determines whether the base orbifold $\mathcal{O}_{M}$ (resp. $\mathcal{O}_{N}$) is orientable, and determines the isomorphism type of the orbifold fundamental group $\pi_{1}(\mathcal{O}_{M})$ (resp. $\pi_{1}(\mathcal{O}_{N})$). Thus, $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$ implies that $\mathcal{O}_{M}$ is non-orientable if and only if $\mathcal{O}_{N}$ is non-orientable. In addition, there is a one-to-one correspondence between the conjugacy classes of maximal torsion subgroups in the orbifold fundamental group with the singular points of a 2-orbifold [^2] We only allow discrete singular set in a 2-orbifold. with boundary. Therefore, $\mathcal{O}_{M}$ contains two singular points of index $2$ if and only if $\pi_{1}(\mathcal{O}_{M})\cong\pi_{1}(\mathcal{O}_{N})$ contains two conjugacy classes of maximal torsion subgroups isomorphic to $\mathbb{Z}/2\mathbb{Z}$, which implies that $\mathcal{O}_{N}$ also contains two singular points of index $2$. ∎
Lemma 5.8.
Let $M$ and $N$ be compact, orientable 3-manifolds with empty or toral boundary such that $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$. Suppose $M$ is aspherical, and $M$ contains an embedded Klein bottle. Then $N$ is also aspherical and contains an embedded Klein bottle.
Proof.
Since $M$ contains an embedded Klein bottle, $M$ is not homeomorphic to $D^{2}\times S^{1}$ or $S^{2}\times S^{1}$. It follows from Corollary 5.3 that $N$ is also aspherical. In particular, $M$ is boundary-incompressible, and it follows from Lemma 5.1 that $N$ is also boundary-incompressible.
It suffices to show that if condition (i) or (ii) in Lemma 5.6 holds for $M$, then the same condition holds for $N$.
First, we suppose that $M$ satisfies condition (i) in Lemma 5.6, i.e. $M$ is a closed $Sol$-manifold which is not fibered over $S^{1}$. Since $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$, it follows from [40, Corollary 4.12] that $N$ is also closed, and [54, Theorem 8.4] shows that $N$ also admits $Sol$-geometry. In addition, [31, Corollary 1.2] shows that the profinite completion of fundamental group of a compact, orientable 3-manifold determines whether this 3-manifold fibers over $S^{1}$. Thus, $N$ is also non-fibered, and satisfies condition (i) of Lemma 5.6.
Next, we suppose that $M$ satisfies condition (ii) in Lemma 5.6. Similarly, [54, Theorem 8.4] implies that $N$ is not a $Sol$-manifold. [56, Theorem B] shows that the profinite completion of fundamental group detects the JSJ-decomposition of a compact, orientable, irreducible 3-manifold with empty or incompressible toral boundary, which does not admit $Sol$-geometry. In fact, [56, Theorem B] was originally proven for closed manifolds, but the proof can be easily generalized to the bounded case, see for example [57, Theorem 5.6]. In particular, there is a one-to-one correspondence between the Seifert fibered pieces of $M$ and $N$, such that the corresponding pieces are profinitely isomorphic. Since one of the Seifert pieces of $M$ has a base orbifold which is either non-orientable or contains two singular points of index $2$, and this Seifert piece is obviously boundary incompressible and aspherical, it follows from Lemma 5.7 that the corresponding Seifert piece of $N$ also has a base orbifold which is either non-orientable or contains two singular points of index $2$. Therefore, $N$ also satisfies condition (ii) in Lemma 5.6. ∎
Remark 5.9.
The aspherical condition is important. In fact, there exist two lens spaces with isomorphic fundamental groups, such that one of them contains a Klein bottle while the other one does not; see [7, Corollary 6.4].
5.4. Profinite detection of exceptional Dehn fillings
Proposition 5.10.
Suppose $M$ and $N$ are orientable finite-volume hyperbolic 3-manifolds with cusps, and $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$. Let $\partial_{i_{0}}M$ be a boundary component of $M$. Then there exists a boundary component $\partial_{j_{0}}N$ of $N$, and an isomorphism $\psi:\pi_{1}\partial_{i_{0}}M\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}\partial_{j_{0}}N$, such that
  1. (1)
    $\overline{\psi}(\mathcal{E}(M,\partial_{i_{0}}M))=\mathcal{E}(N,\partial_{j_{0}}N)$, where $\overline{\psi}$ is the bijection $\text{slp}(\partial_{i_{0}}M)\to\text{slp}(\partial_{j_{0}}N)$;
  2. (2)
    $\overline{\psi}(\mathcal{E}_{s}(M,\partial_{i_{0}}M))=\mathcal{E}_{s}(N,\partial_{j_{0}}N)$;
  3. (3)
    $\overline{\psi}(\mathcal{E}_{t}(M,\partial_{i_{0}}M))=\mathcal{E}_{t}(N,\partial_{j_{0}}N)$;
  4. (4)
    $\overline{\psi}(\mathcal{E}_{k}(M,\partial_{i_{0}}M))=\mathcal{E}_{k}(N,\partial_{j_{0}}N)$.
Proof.
According to Theorem A, there is a boundary component $\partial_{j_{0}}N$ of $N$, and an isomorphism $\psi:\pi_{1}\partial_{i_{0}}M\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}\partial_{j_{0}}N$, such that for any $c\in\text{slp}(\partial_{i_{0}}M)$, $\widehat{\pi_{1}M_{c}}\cong\widehat{\pi_{1}N_{\overline{\psi}(c)}}$ by an isomorphism which respects the peripheral structure.
According to Proposition 2.2, $M_{c}$ is finite-volume hyperbolic if and only if $N_{\overline{\psi}(c)}$ is finite-volume hyperbolic. Hence, by definition of exceptional Dehn filling, $c\in\mathcal{E}(M,\partial_{i_{0}}M)$ if and only if $\overline{\psi}(c)\in\mathcal{E}(N,\partial_{j_{0}}N)$, which proves (1).
Since $M_{c}$ and $N_{\overline{\psi}(c)}$ have the same number of boundary components, Corollary 5.3 implies that $M_{c}$ is spherical if and only if $N_{\overline{\psi}(c)}$ is aspherical, which proves (2).
Similarly, (3) follows from Lemma 5.4 since the isomorphism $\widehat{\pi_{1}M_{c}}\cong\widehat{\pi_{1}N_{\overline{\psi}(c)}}$ respects the peripheral structure, and (4) follows from Lemma 5.8. ∎
As a typical example, when $M=X_{K}$ and $N=X_{K^{\prime}}$ are hyperbolic knot complements, we have established a standard parametrization $\text{slp}(\partial M)=\mathbb{Q}\cup\{\infty\}$ and $\text{slp}(\partial N)=\mathbb{Q}\cup\{\infty\}$ in Section 4. Thus, $\mathcal{E}(M,\partial M)$ and $\mathcal{E}(N,\partial N)$ can be viewed as subsets of ${}\mathbb{Q}\cup\{\infty\}$. For brevity of notation, when $M=X_{K}$, we simply denote $\mathcal{E}(M,\partial M)$ as $\mathcal{E}(K)$.
Based on Theorem 4.5, we obtain a more precise version of Proposition 5.10 for hyperbolic knot complements.
Proposition 5.11.
Suppose $K$ and $K^{\prime}$ are hyperbolic knots in $S^{3}$, and $\widehat{\pi_{1}X_{K}}\cong\widehat{\pi_{1}X_{K^{\prime}}}$. Then there exists $\sigma\in\{1,-1\}$, such that
  1. (1)
    $\mathcal{E}(K)=\sigma\cdot\mathcal{E}(K^{\prime})$;
  2. (2)
    $\mathcal{E}_{s}(K)=\sigma\cdot\mathcal{E}_{s}(K^{\prime})$;
  3. (3)
    $\mathcal{E}_{t}(K)=\sigma\cdot\mathcal{E}_{t}(K^{\prime})$;
  4. (4)
    $\mathcal{E}_{k}(K)=\sigma\cdot\mathcal{E}_{k}(K^{\prime})$.
Proof.
As in the proof of Proposition 5.10, let $\psi:\pi_{1}\partial X_{K}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}\partial X_{K^{\prime}}$ denote the isomorphism such that $\widehat{\pi_{1}(X_{K})_{c}}\cong\widehat{\pi_{1}(X_{K^{\prime}})_{\overline{\psi}(c)}}$ for any $c\in\text{slp}^{\ast}(\partial X_{K})$. According to Theorem 4.5, under the parametrization by ${}\mathbb{Q}\cup\{\infty\}$, the map $\overline{\psi}:\text{slp}(\partial X_{K})=\mathbb{Q}\cup\{\infty\}\to\text{slp}(\partial X_{K^{\prime}})=\mathbb{Q}\cup\{\infty\}$ is a multiplication by $\sigma\in\{\pm 1\}$. Therefore, the results of this proposition follow directly from Proposition 5.10. ∎
6. Profinite rigidity from exceptional slopes
In this section, we shall prove Theorem B. The key concept we use in this section is the geometric intersection number. For two slopes on a torus $c_{1},c_{2}\in\text{slp}(T^{2})$, we denote by $\Delta(c_{1},c_{2})$ the geometric intersection number of $c_{1}$ and $c_{2}$. Note that $\Delta(c_{1},c_{2})$ is a non-negative integer, and $\Delta(c_{1},c_{2})=0$ if and only if $c_{1}=c_{2}$.
Example 6.1.
Let $K$ be a knot in $S^{3}$, and we parametrize $\text{slp}(\partial X_{K})$ by ${}\mathbb{Q}\cup\{\infty\}$ as in Section 4. For $r_{1},r_{2}\in\mathbb{Q}\cup\{\infty\}$, we can express them as reduced fractions $r_{1}=\frac{p_{1}}{q_{1}}$ and $r_{2}=\frac{p_{2}}{q_{2}}$. Then $\Delta(r_{1},r_{2})=\left|\det\left(\begin{smallmatrix}p_{1}&q_{1}\\ p_{2}&q_{2}\end{smallmatrix}\right)\right|$.
The next lemma is an obvious but useful observation.
Lemma 6.2.
Let $T_{1}^{2}$ and $T_{2}^{2}$ denote two tori, and let $\psi:\pi_{1}(T_{1}^{2})\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}(T_{2}^{2})$ be a group isomorphism. Then, for any slopes $c_{1},c_{2}\in\text{slp}(T_{1}^{2})$, $\Delta(c_{1},c_{2})=\Delta(\overline{\psi}(c_{1}),\overline{\psi}(c_{2}))$.
Proof.
One can view $\psi:\pi_{1}(T_{1}^{2})\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}(T_{2}^{2})$ as a homeomorphism $\Psi:T_{1}^{2}\to T_{2}^{2}$, and directly derive this lemma by definition. ∎
6.1. Some Dehn-surgeries on the Whitehead link
We start from the easiest example, so as to illustrate how our method works. In fact, the manifolds listed in Theorem B can be characterised by their extremely special patterns of exceptional Dehn fillings.
For $r\in\mathbb{Q}\cup\{\infty\}$, let $\mathcal{W}(r)$ denote the 3-manifold obtained by $r$-surgery on one component of the Whitehead link (Figure 1(a)).
Proposition 6.3 ([22, Theorem 1.1] ).
Suppose $N$ is a finite-volume hyperbolic 3-manifold with one cusp.
  1. (1)
    $\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{t}(N,\partial N)\}=6$ if and only if $N$ is homeomorphic to $\mathcal{W}(-2)$.
  2. (2)
    $\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{t}(N,\partial N)\}=7$ if and only if $N$ is homeomorphic to $\mathcal{W}(\frac{5}{2})$.
  3. (3)
    $\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{t}(N,\partial N)\}=8$ if and only if $N$ is homeomorphic to either $\mathcal{W}(-1)$ (the figure-eight knot complement) or $\mathcal{W}(5)$ (the figure-eight sibling manifold).
Proposition 6.4 ([33, Corollary 1.5] ).
Suppose $N$ is a finite-volume hyperbolic 3-manifold with one cusp. Then $\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{k}(N,\partial N)\}=5$ if and only if $N$ is homeomorphic to $\mathcal{W}(-4)$.
Proof.
The “if” direction follows from [37, Table A.4], and the “only if” direction follows from [33, Corollary 1.5]. ∎
It is well-known that all the exceptional slopes on the Whitehead link (Figure 1(a)) are $\infty,0,1,2,3,4$ [37]. Thus, the above listed $\mathcal{W}(-2)$, $\mathcal{W}(\frac{5}{2})$, $\mathcal{W}(-1)$, $\mathcal{W}(5)$ and $\mathcal{W}(-4)$ are all finite-volume hyperbolic 3-manifolds with one cusp. Note that the Dehn surgery coefficients in [22] differ from the ones we use here by a $\pm$-sign, due to the chirality of the Whitehead link.
Theorem 6.5.
The manifolds $\mathcal{W}(-2)$, $\mathcal{W}(\frac{5}{2})$, $\mathcal{W}(-4)$, $\mathcal{W}(-1)$ and $\mathcal{W}(5)$ are all profinitely rigid among compact orientable 3-manifolds.
Proof.
We start from the example of $\mathcal{W}(-2)$. Let $N$ be a compact orientable 3-manifold so that $\widehat{\pi_{1}\mathcal{W}(-2)}\cong\widehat{\pi_{1}N}$. Then Proposition 2.2 implies that $N$ is also finite-volume hyperbolic with exactly one cusp. Therefore, Proposition 5.10 implies that there is an isomorphism $\psi:\pi_{1}(\partial\mathcal{W}(-2))\to\pi_{1}(\partial N)$, such that $\overline{\psi}(\mathcal{E}_{t}(\mathcal{W}(-2),\partial\mathcal{W}(-2)))=\mathcal{E}_{t}(N,\partial N)$. Note that $\overline{\psi}$ is induced by a group isomorphism, so Lemma 6.2 implies that $\overline{\psi}$ preserves the geometric intersection number $\Delta(\cdot,\cdot)$. Thus, $\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{t}(N,\partial N)\}=\max\{\Delta(c_{1}^{\prime},c_{2}^{\prime})\mid c_{1}^{\prime},c_{2}^{\prime}\in\mathcal{E}_{t}(\mathcal{W}(-2),\partial\mathcal{W}(-2))\}=6$. According to Proposition 6.3 (1), the only possibility is $N\cong\mathcal{W}(-2)$.
The profinite rigidity of $\mathcal{W}(\frac{5}{2})$ and $\mathcal{W}(-4)$ follow from the same proof as $\mathcal{W}(-2)$, based on Proposition 6.3 (2) and Proposition 6.4 respectively.
Moving on to $\mathcal{W}(-1)$ and $\mathcal{W}(5)$, the same proof implies that if a compact orientable 3-manifold $N$ is profinitely isomorphic to $\mathcal{W}(-1)$ or $\mathcal{W}(5)$, then $N$ is homeomorphic to one of $\mathcal{W}(-1)$ and $\mathcal{W}(5)$. Thus, it suffices to show that $\mathcal{W}(-1)$ and $\mathcal{W}(5)$ are not profinitely isomorphic. In fact, $\mathcal{W}(-1)$ is the figure-eight knot complement, while the figure-eight sibling manifold $\mathcal{W}(5)$ is not a knot complement. Therefore, Theorem 4.1 guarantees that $\widehat{\pi_{1}\mathcal{W}(-1)}\not\cong\widehat{\pi_{1}\mathcal{W}(5)}$, finishing the proof. ∎
6.2. Whitehead link, Whitehead sister link, $\frac{3}{10}$ two-bridge link, and $M_{14}$
Recall that $\mathcal{W}$, $\mathcal{L}$ and $\mathcal{WS}$ denote the Whitehead link, the two-bridge link associated to the rational number $\frac{3}{10}$, and the Whitehead sister link (i.e. the $(-2,3,8)$-Pretzel link) respectively. $M_{14}$ is the double branched cover along the tangle $\mathcal{Q}$ in $S^{2}\times I$ (Figure 6) defined in [24].
Proposition 6.6 ([24, Corollary 1.3] ).
Suppose $N$ is a finite-volume hyperbolic 3-manifold with two cusps, and let $\partial_{0}N$ be a boundary component of $N$.
  1. (1)
    $\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{t}(N,\partial_{0}N)\}=4$ if and only if $N$ is homeomorphic to one of $S^{3}\setminus\mathring{N}(\mathcal{W})$, $S^{3}\setminus\mathring{N}(\mathcal{L})$, and $M_{14}$.
  2. (2)
    $\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{t}(N,\partial_{0}N)\}=5$ if and only if $N$ is homeomorphic to $S^{3}\setminus\mathring{N}(\mathcal{WS})$.
For simplicity, we denote $M_{1}=S^{3}\setminus\mathring{N}(\mathcal{W})$, $M_{2}=S^{3}\setminus\mathring{N}(\mathcal{L})$, and $M_{3}=S^{3}\setminus\mathring{N}(\mathcal{WS})$ following the notation of [24]. It is well-known that $M_{1}$, $M_{2}$ and $M_{3}$ are finite-volume hyperbolic 3-manifolds, and [24, Lemma 23.13] proved that $M_{14}$ is also finite-volume hyperbolic.
Lemma 6.7.
Any two of $\widehat{\pi_{1}M_{1}}$, $\widehat{\pi_{1}M_{2}}$, $\widehat{\pi_{1}M_{14}}$ are not isomorphic.
Proof.
We first show that $\widehat{\pi_{1}M_{1}}\not\cong\widehat{\pi_{1}M_{14}}$ and $\widehat{\pi_{1}M_{2}}\not\cong\widehat{\pi_{1}M_{14}}$. In fact, $M_{1}$ and $M_{2}$ are both hyperbolic link complements in $S^{3}$, while in the proof of [24, Lemma 24.2], Gordon-Wu showed that $M_{14}$ is not a link complement. Therefore, Theorem 4.1 implies $\widehat{\pi_{1}M_{1}}\not\cong\widehat{\pi_{1}M_{14}}$ and $\widehat{\pi_{1}M_{2}}\not\cong\widehat{\pi_{1}M_{14}}$.
Next, we show that $\widehat{\pi_{1}M_{1}}\not\cong\widehat{\pi_{1}M_{2}}$. Suppose on the contrary that $f:\widehat{\pi_{1}M_{1}}\to\widehat{\pi_{1}M_{2}}$ is an isomorphism. Fix the conjugacy representatives of peripheral subgroups $\pi_{1}\partial_{i}M_{1}\subseteq\pi_{1}M_{1}$ and $\pi_{1}\partial_{i}M_{2}\subseteq\pi_{1}M_{2}$ ($i=1,2$). Then Theorem 3.2 (1) implies that there exists $i\in\{1,2\}$ such that $f(\overline{\pi_{1}\partial_{1}M_{1}})$ is a conjugate of $\overline{\pi_{1}\partial_{i}M_{2}}$ in $\widehat{\pi_{1}M_{2}}$. By composing with a conjugation, we may simply assume $f(\overline{\pi_{1}\partial_{1}M_{1}})=\overline{\pi_{1}\partial_{i}M_{2}}$. Recall that $\overline{\pi_{1}\partial_{i}M_{j}}\cong\widehat{\pi_{1}\partial_{i}M_{j}}\cong\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}\pi_{1}\partial_{i}M_{j}$ by Proposition 2.3. Then the abelianization of profinite groups [57, Proposition 2.11] yields the following commutative diagram.
[…]
In fact, the vertical maps in the commutative diagram are induced by the inclusions $\rho:\partial_{1}M_{1}\to M_{1}$ and $\iota:\partial_{i}M_{2}\to M_{2}$, i.e. ^$\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}H_{1}(\partial_{1}M_{1};\mathbb{Z})\xrightarrow{1\otimes\rho_{\ast}}\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}H_{1}(M_{1};\mathbb{Z})$ and likewise.
Note that the linking number of the two components of $\mathcal{W}$ is 00, so each boundary component of $M_{1}=S^{3}\setminus\mathring{N}(\mathcal{W})$ contains a null-homologous boundary slope. In other words, $\rho_{\ast}:H_{1}(\partial_{1}M_{1};\mathbb{Z})\to H_{1}(M_{1};\mathbb{Z})$ is not injective. Note that $H_{1}(\partial_{1}M_{1};\mathbb{Z})$ is a finitely generated Abelian group, so $H_{1}(\partial_{1}M_{1};\mathbb{Z})$ injects into ^$\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}H_{1}(\partial_{1}M_{1};\mathbb{Z})$, and ^$\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}H_{1}(\partial_{1}M_{1};\mathbb{Z})\xrightarrow{1\otimes\rho_{\ast}}\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}H_{1}(M_{1};\mathbb{Z})$ is not injective. On the other hand, the linking number of the two components of $\mathcal{L}$ is ±$\pm 3$ (depending on the orientation). Thus, the two boundary components of $M_{2}=S^{3}\setminus\mathring{N}(\mathcal{L})$ are both $H_{1}$-injective, i.e. $\iota_{\ast}:H_{1}(\partial_{i}M_{2};\mathbb{Z})\to H_{1}(M_{2};\mathbb{Z})$ is injective. Since $\widehat{\mathbb{Z}}$ is torsion-free, and hence a flat $\mathbb{Z}$-module, the homomorphism ^$\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}H_{1}(\partial_{i}M_{2};\mathbb{Z})\xrightarrow{1\otimes\iota_{\ast}}\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}H_{1}(M_{2};\mathbb{Z})$ is injective. Therefore, the following diagram shall not commute, which yields a contradiction.
[…]
Theorem 6.8.
$M_{1},M_{2},M_{3},M_{14}$ are profinitely rigid among compact, orientable 3-manifolds.
Proof.
Recall that $M_{1},M_{2},M_{3},M_{14}$ are finite-volume hyperbolic 3-manifolds with two cusps. According to Proposition 2.2, any compact orientable 3-manifold profinitely isomorphic to $M_{1}$, $M_{2}$, $M_{3}$ or $M_{14}$ is also a finite-volume hyperbolic 3-manifolds with two cusps.
We first prove the profinite rigidity of $M_{3}$. Suppose $N$ is a finite-volume hyperbolic 3-manifold with two cusps such that $\widehat{\pi_{1}M_{3}}\cong\widehat{\pi_{1}N}$. Let $\partial_{0}M_{3}$ be a boundary component of $M_{3}$ such that $\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{t}(M_{3},\partial_{0}M_{3})\}=5$ according to Proposition 6.6. Then similar to the proof of Theorem 6.5, we can derive from Proposition 5.10 and Lemma 6.2 that there exists a boundary component $\partial_{i}N$ of $N$ such that $\max\{\Delta(c_{1}^{\prime},c_{2}^{\prime})\mid c_{1}^{\prime},c_{2}^{\prime}\in\mathcal{E}_{t}(N,\partial_{i}N)\}=\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{t}(M_{3},\partial_{0}M_{3})\}=5$. Proposition 6.6 (2) then implies that $N\cong M_{3}$.
Next, we show the profinite rigidity of $M_{1}$, $M_{2}$ and $M_{14}$. Suppose $j\in\{1,2,14\}$ and $N$ is a finite-volume hyperbolic 3-manifold with two cusps such that $\widehat{\pi_{1}M_{j}}\cong\widehat{\pi_{1}N}$. Similar to the proof for $M_{3}$, there exists a boundary component $\partial_{i}N$ of $N$ such that $\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{t}(N,\partial_{i}N)\}=4$. Proposition 6.6 (1) then implies that $N$ is homeomorphic to one of $M_{1}$, $M_{2}$ and $M_{14}$. We have shown in Lemma 6.7 that any two of $M_{1}$, $M_{2}$ and $M_{14}$ are not profinitely isomorphic. Hence, the only remaining possibility is $N\cong M_{j}$. ∎
6.3. More hyperbolic knot complements in $S^{3}$
In this subsection, we consider three families of hyperbolic knots.
  1. (1)
    The twist knots $\mathcal{K}_{n}$ (Figure 2) can also be viewed as a $\frac{1}{n}$-Dehn surgery on one component of the Whitehead link (Figure 1(a)). For example, $\mathcal{K}_{0}$ is the unknot, $\mathcal{K}_{1}$ is the trefoil, and $\mathcal{K}_{-1}$ is the figure-eight knot. In fact, according to [37], for all $n\in\mathbb{Z}\setminus\{0,1\}$, $\mathcal{K}_{n}$ is a hyperbolic knot.
  2. (2)
    The braid knots $\mathcal{J}_{n}$ represented by $\sigma_{2}^{-1}\sigma_{1}(\sigma_{1}\sigma_{2})^{3n}$ (Figure 3) can be viewed as a $\frac{1}{n}$-surgery on one component of the two-bridge link $\mathcal{L}$ shown in Figure 1(c). Again, [37] shows that $\mathcal{J}_{n}$ is hyperbolic if and only if $n\in\mathbb{Z}\setminus\{-1,0,1\}$.
  3. (3)
    The Eudave-Muñoz knots $K(3,1,n,0)$ defined in [18] (Figure 4(a)) can be viewed as a $-\frac{1}{n}$-surgery on the unknotted component of the Whitehead sister link $\mathcal{WS}$ shown in Figure 1(b). Indeed, [18, Proposition 2.2] shows that $K(3,1,n,0)$ is hyperbolic for all $n\in\mathbb{Z}\setminus\{0\}$, and the mirror-image of $K(3,1,n,0)$ is $K(2,-1,1-n,0)$ by [18, Proposition 1.4].
Proposition 6.9 ([24, Theorem 24.4] ).
Suppose ${K}$ is a hyperbolic knot in $S^{3}$.
  1. (1)
    If there exist $c_{1},c_{2}\in\mathcal{E}_{t}(K)$ such that $\Delta(c_{1},c_{2})=8$, then $K$ is the figure-eight knot $\mathcal{K}_{-1}={4_{1}}$.
  2. (2)
    If there exist $c_{1},c_{2}\in\mathcal{E}_{t}(K)$ such that $\Delta(c_{1},c_{2})=4$, then $K$ is equivalent or mirror image to $\mathcal{K}_{n}$ ($n\in\mathbb{Z}\setminus\{0,1\}$) or $\mathcal{J}_{n}$ ($n\in\mathbb{Z}\setminus\{-1,0,1\}$).
  3. (3)
    If there exist $c_{1},c_{2}\in\mathcal{E}_{t}(K)$ such that $\Delta(c_{1},c_{2})=5$, then $K$ is the Eudave-Muñoz knot $K(3,1,n,0)$ or its mirror-image $K(2,-1,1-n,0)$ for some $n\in\mathbb{Z}\setminus\{0\}$.
Proposition 6.10.
  1. (1)
    $\mathcal{E}_{t}(\mathcal{K}_{-1})=\{-4,0,4\}$. In particular, $\Delta(-4,4)=8$
  2. (2)
    For $n\in\mathbb{Z}\setminus\{-1,0,1\}$, $\mathcal{E}_{t}(\mathcal{K}_{n})=\{0,4\}$, and $\Delta(0,4)=4$.
  3. (3)
    For $n\in\mathbb{Z}\setminus\{-1,0,1\}$, $\mathcal{E}_{t}(\mathcal{J}_{n})=\{2-9n,-2-9n\}$, and $\Delta(2-9n,-2-9n)=4$.
  4. (4)
    For $n\in\mathbb{Z}\setminus\{0\}$, $\{25n-9,25n-\frac{13}{2}\}\subseteq\mathcal{E}_{t}(K(3,1,n,0))$, and $\Delta(25n-9,25n-\frac{13}{2})=5$. In addition, $25n-\frac{13}{2}$ is the unique half-integral element in $\mathcal{E}_{t}(K(3,1,n,0))$.
Proof.
(1) and (2) are exactly (4) and (3) of [12, Theorem 1.1], and the calculation of geometric intersection number follows from Example 6.1. (3) is proven in [24, Corollary 24.5], see also [37].
Now we move on to (4). The fact that $\{25n-9,25n-\frac{13}{2}\}\subseteq\mathcal{E}_{t}(K(3,1,n,0))$ follows from [19, Proposition 5.4 (1) and (11)] [^3] In fact, there is a calculation error in the Seifert invariants in [19, Proposition 5.4 (1)], but the surgery coefficient is still correct., together with the fact that $K(3,1,n,0)$ is the mirror-image of $K(2,-1,1-n,0)$ [18, Proposition 1.4]. In particular, when $n=1$, $K(3,1,n,0)$ is the $(-2,3,7)$-Pretzel knot, and $\mathcal{E}_{t}(K(3,1,n,0))=\{16,\frac{37}{2},20\}$ strictly contains $\{25n-9,25n-\frac{13}{2}\}=\{16,\frac{37}{2}\}$. Nevertheless, any hyperbolic knot in $S^{3}$ has at most one non-integral toroidal slope as shown by [25, Theorem 1]. Thus, for any $n\in\mathbb{Z}\setminus\{0\}$, $\mathcal{E}_{t}(K(3,1,n,0))$ contains exactly one half-integral element, which is $25n-\frac{13}{2}$ stated above. ∎
Lemma 6.11.
  1. (1)
    For any $m\in\mathbb{Z}\setminus\{0,1\}$ and $n\in\mathbb{Z}\setminus\{-1,0,1\}$, $\widehat{\pi_{1}X_{\mathcal{K}_{m}}}\not\cong\widehat{\pi_{1}X_{\mathcal{J}_{n}}}$.
  2. (2)
    For $m,n\in\mathbb{Z}\setminus\{-1,0,1\}$, $\widehat{\pi_{1}X_{\mathcal{J}_{m}}}\cong\widehat{\pi_{1}X_{\mathcal{J}_{n}}}$ if and only if $m=\pm n$; in this case, $\mathcal{J}_{m}$ and $\mathcal{J}_{n}$ are either equivalent or the mirror-image of each other.
  3. (3)
    For $m,n\in\mathbb{Z}\setminus\{0\}$, $\widehat{\pi_{1}X_{K(3,1,m,0)}}\cong\widehat{\pi_{1}X_{K(3,1,n,0)}}$ if and only if $m=n$.
  4. (4)
    For $m,n\in\mathbb{Z}\setminus\{0,1\}$, $\widehat{\pi_{1}X_{\mathcal{K}_{m}}}\cong\widehat{\pi_{1}X_{\mathcal{K}_{n}}}$ if and only if $m=n$.
Proof.
Recall by Proposition 5.11 that if two hyperbolic knot complements $X_{K}$ and $X_{K^{\prime}}$ are profinitely isomorphic, then either $\mathcal{E}_{t}(K)=\mathcal{E}_{t}(K^{\prime})$ or $\mathcal{E}_{t}(K)=-\mathcal{E}_{t}(K^{\prime})$.
According to Proposition 6.10, $0\in\mathcal{E}_{t}(\mathcal{K}_{m})$ while $0\notin\mathcal{E}_{t}(\mathcal{J}_{n})$. Thus, this implies that $\widehat{X_{\mathcal{K}_{m}}}\not\cong\widehat{X_{\mathcal{J}_{n}}}$, which proves (1).
Similarly, if $\widehat{\pi_{1}X_{\mathcal{J}_{m}}}\cong\widehat{\pi_{1}X_{\mathcal{J}_{n}}}$, then $\{-2-9m,2-9m\}=\mathcal{E}_{t}(K)=\pm\mathcal{E}_{t}(K^{\prime})=\pm\{-2-9n,2-9n\}$ by Proposition 6.10, which implies $m=\pm n$. In addition, $\mathcal{J}_{n}$ can be viewed as a $\frac{1}{n}$-surgery on the link $\mathcal{L}$, while $\mathcal{L}$ is achiral, i.e. equivalent to its mirror image. Thus, the $\frac{1}{n}$-surgery on $\mathcal{L}$ yields the mirror-image of the $-\frac{1}{n}$-surgery on $\mathcal{L}$. This implies that $\mathcal{J}_{n}$ and $\mathcal{J}_{-n}$ are mirror-image of each other. In particular, $X_{\mathcal{J}_{n}}$ is homeomorphic to $X_{\mathcal{J}_{-n}}$, so $\widehat{\pi_{1}X_{\mathcal{J}_{n}}}\cong\widehat{\pi_{1}X_{\mathcal{J}_{-n}}}$. This finishes the proof of (2).
For (3), note that $25n-\frac{13}{2}$ is the unique half-integral element in $\mathcal{E}_{t}(K(3,1,n,0))$. Thus, if $\widehat{\pi_{1}X_{K(3,1,m,0)}}\cong\widehat{\pi_{1}X_{K(3,1,n,0)}}$, the same reasoning implies $25m-\frac{13}{2}=\pm(25n-\frac{13}{2})$. Since $m$ and $n$ are integers, the only possibility is $m=n$.
(4) is difficult to detect from the toroidal slopes; instead, we can detect these knots through the profinite rigidity of Alexander polynomials [46]. In fact, Ueki [46, Theorem 1.1] shows that if two knot complements $X_{K}$ and $X_{K^{\prime}}$ are profinitely isomorphic, then their Alexander polynomials $\Delta_{K}(t)$ and $\Delta_{K^{\prime}}(t)$ coincide up to a multiplicative unit in $\mathbb{Z}[t^{\pm}]$. According to [42, Chapter 7.B, Exercise 7], the Alexander polynomial of the twist knot $\mathcal{K}_{m}$ is $\Delta_{\mathcal{K}_{m}}(t)\doteq mt^{2}+(1-2m)t+m$. It is easy to verify that for any $m\neq n\in\mathbb{Z}$, $\Delta_{\mathcal{K}_{m}}(t)$ and $\Delta_{\mathcal{K}_{n}}(t)$ are not equivalent up to multiplicative unit, and therefore, $\widehat{\pi_{1}X_{\mathcal{K}_{m}}}\not\cong\widehat{\pi_{1}X_{\mathcal{K}_{n}}}$. ∎
Theorem 6.12.
The knot complements $X_{\mathcal{K}_{n}}$ ($n\in\mathbb{Z}\setminus\{0,1\}$), $X_{\mathcal{J}_{n}}$ ($n\in\mathbb{Z}\setminus\{-1,0,1\}$), and $X_{K(3,1,-n,0)}$ ($n\in\mathbb{Z}\setminus\{0\}$) are profinitely rigid among all compact orientable 3-manifolds.
Proof.
For brevity of notation, let $K$ denote one of the above listed knots, i.e. $K$ belongs to $\mathcal{K}_{n}$ ($n\in\mathbb{Z}\setminus\{0,1\}$), ${\mathcal{J}_{n}}$ ($n\in\mathbb{Z}\setminus\{-1,0,1\}$), and ${K(3,1,-n,0)}$ ($n\in\mathbb{Z}\setminus\{0\}$), and we denote $M=X_{K}$.
Suppose $N$ is a compact orientable 3-manifold such that $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$. Then according to Theorem 4.1, $N\cong X_{K^{\prime}}$ is also a hyperbolic knot complement. In addition, $\mathcal{E}_{t}(K^{\prime})=\pm\mathcal{E}_{t}(K)$ according to Proposition 5.11. According to Example 6.1, the $\pm$-sign does not affect the geometric intersection number $\Delta(\cdot,\cdot)$, so the possible geometric intersection numbers of toroidal slopes of $K$ and $K^{\prime}$ coincide.
When $K=\mathcal{K}_{-1}$, Proposition 6.10 (1) implies that there exist $c_{1},c_{2}\in\mathcal{E}_{t}(K)$ such that $\Delta(c_{1},c_{2})=8$. By the above reasoning, there exist $c_{1}^{\prime},c_{2}^{\prime}\in\mathcal{E}_{t}(K^{\prime})$ such that $\Delta(c_{1}^{\prime},c_{2}^{\prime})=8$, and $K^{\prime}$ is equivalent to the figure-eight knot $K_{-1}$ according to Proposition 6.9 (1). Thus, $N\cong M$ is the figure-eight knot complement.
When $K=K(3,1,n,0)$, Proposition 6.10 (4) implies that there exist $c_{1},c_{2}\in\mathcal{E}_{t}(K)$ such that $\Delta(c_{1},c_{2})=5$. The above reasoning shows that there exist $c_{1}^{\prime},c_{2}^{\prime}\in\mathcal{E}_{t}(K^{\prime})$ such that $\Delta(c_{1}^{\prime},c_{2}^{\prime})=5$. According to Proposition 6.9 (3), $K^{\prime}$ is either equivalent or mirror image to a Eudave-Muñoz knot $K(3,1,n^{\prime},0)$ ($n^{\prime}\in\mathbb{Z}\setminus\{0\}$). Thus, $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$ implies $\widehat{\pi_{1}X_{K(3,1,n,0)}}\cong\widehat{\pi_{1}X_{K(3,1,n^{\prime},0)}}$. Lemma 6.11 then guarantees $n=n^{\prime}$, i.e. $N\cong X_{K^{\prime}}\cong X_{K(3,1,n,0)}\cong M$.
When $K=\mathcal{K}_{n}$ or $K=\mathcal{J}_{n}$ for some $n\in\mathbb{Z}\setminus\{-1,0,1\}$, Proposition 6.10 (2) and (3) imply that there exist $c_{1},c_{2}\in\mathcal{E}_{t}(K)$ such that $\Delta(c_{1},c_{2})=4$. As a result, there exist $c_{1}^{\prime},c_{2}^{\prime}\in\mathcal{E}_{t}(K^{\prime})$ such that $\Delta(c_{1}^{\prime},c_{2}^{\prime})=4$, and Proposition 6.9 (2) implies that $K^{\prime}$ is either equivalent or mirror image to $\mathcal{K}_{n^{\prime}}$ ($n^{\prime}\in\mathbb{Z}\setminus\{0,1\}$) or $\mathcal{J}_{n^{\prime}}$ ($n^{\prime}\in\mathbb{Z}\setminus\{-1,0,1\}$). In addition, $\widehat{\pi_{1}X_{K}}\cong\widehat{\pi_{1}X_{K^{\prime}}}$. According to (1), (2), (4) of Lemma 6.11, these candidates can be profinitely distinguished from each other, which ensures that $K^{\prime}$ can only be equivalent or mirror image to $K$. Thus, $N\cong X_{K^{\prime}}\cong X_{K}=M$, finishing the proof. ∎
6.4. The Berge manifold
The complement of the link $L_{A}$ shown in Figure 5(a) can also be viewed as the complement of the $W_{3}^{1}W_{7}^{-3}$ braid knot in $D^{2}\times S^{1}$, which is conventionally called the Berge manifold. Indeed, this manifold is hyperbolic according to [6, Theorem 3.2]. In addition, we have the following characterization.
Proposition 6.13 ([6, Corollary 2.9] ).
Suppose $N$ is a finite-volume hyperbolic 3-manifold with two cusps. There exist three different slopes $c_{1},c_{2},c_{3}$ on one boundary component of $N$ such that $N_{c_{1}}\cong N_{c_{2}}\cong N_{c_{3}}\cong D^{2}\times S^{1}$ if and only if $N$ is homeomorphic to $S^{3}\setminus\mathring{N}(L_{A})$.
Theorem 6.14.
The Berge manifold $S^{3}\setminus\mathring{N}(L_{A})$ is profinitely rigid among all compact orientable 3-manifolds.
Proof.
Suppose $N$ is a compact orientable 3-manifold such that $\widehat{\pi_{1}N}\cong\widehat{\pi_{1}(S^{3}\setminus L_{A})}$. Since $S^{3}\setminus\mathring{N}(L_{A})$ is a finite-volume hyperbolic 3-manifold according to [6, Theorem 3.2], Proposition 2.2 implies that $N$ is also a finite-volume hyperbolic 3-manifold with two cusps.
Proposition 6.13 shows that there are three different slopes on one boundary component of $S^{3}\setminus\mathring{N}(L_{A})$ such that the corresponding Dehn filling is homeomorphic to $D^{2}\times S^{1}$. According to Theorem A, there exist three different slopes $c_{1},c_{2},c_{3}$ on one boundary component $\partial_{0}N$ of $N$ such that $\widehat{\pi_{1}N_{c_{1}}}\cong\widehat{\pi_{1}N_{c_{2}}}\cong\widehat{\pi_{1}N_{c_{3}}}\cong\widehat{\pi_{1}(D^{2}\times S^{1})}\cong\widehat{\mathbb{Z}}$. Note that the group $\mathbb{Z}$ is profinitely rigid among all finitely generated, residually finite groups. So according to Proposition 2.1, we derive that $\pi_{1}N_{c_{1}}\cong\pi_{1}N_{c_{2}}\cong\pi_{1}N_{c_{3}}\cong\mathbb{Z}$. For each $i=1,2,3$, $N_{c_{i}}$ is a compact orientable 3-manifold whose boundary is a torus, and the only such manifold with infinite cyclic fundamental group is $D^{2}\times S^{1}$. Thus, $N$ admits three different Dehn fillings on $\partial_{0}N$ yielding $D^{2}\times S^{1}$, and Proposition 6.13 implies that $N\cong S^{3}\setminus\mathring{N}(L_{A})$. ∎
6.5. Detecting Klein bottles and Dehn surgeries on $\mathcal{L}$ and $L_{B}$
We first clarify the notations. Recall that $L_{B}$ is the link shown in Figure 5(b). For $u,r,s\in\mathbb{Q}\cup\{\infty\}$, let $L_{B}(u)$ denote $u$-surgery on the $L_{B}^{(1)}$ component of $L_{B}$, while the other two components of $L_{B}$ yield two boundary components of the 3-manifold $L_{B}(u)$; let $L_{B}(u,r)$ denote $u$-surgery on the $L_{B}^{(1)}$ component and $r$-surgery on the $L_{B}^{(2)}$ component, which is a 3-manifold with one torus boundary; let $L_{B}(u,r,s)$ denote $u$-surgery on the $L_{B}^{(1)}$ component, $r$-surgery on the $L_{B}^{(2)}$ component, and $s$-surgery on the $L_{B}^{(3)}$ component, which is a closed 3-manifold.
Note that there is an orientation-preserving homeomorphism of $(S^{3},L_{B})$ which interchanges the $L_{B}^{(2)}$ and $L_{B}^{(3)}$ components. Thus,
(6.15) $L_{B}(u,r,s)\cong L_{B}(u,s,r).$
Recall that $\mathcal{L}$ is the link shown in Figure 1(c). Similarly, let $\mathcal{L}(r)$ denote $r$-surgery on one component of $\mathcal{L}$, and $\mathcal{L}(r,s)$ denote $r$-surgery on one component and $s$-surgery on the other component of $\mathcal{L}$. Note that $\mathcal{L}$ is achiral, i.e. equivalent to its mirror image. Thus,
(6.16) $\mathcal{L}(r)\cong\mathcal{L}(-r)\quad\text{and}\quad\mathcal{L}(r,s)\cong\mathcal{L}(-r,-s).$
Some of the Dehn surgeries on $\mathcal{L}$ and $L_{B}$ can be characterized by their exceptional Dehn fillings, as shown in the following proposition.
Proposition 6.17 ([32, Theorem 1.1] ).
Let $N$ be a cusped finite-volume hyperbolic 3-manifold, and let $\partial_{0}N$ be a boundary component of $N$. Suppose there exist $c_{1},c_{2}\in\text{slp}(\partial_{0}N)$ with $\Delta(c_{1},c_{2})=4$, such that both $N_{c_{1}}$ and $N_{c_{2}}$ contain an embedded Klein bottle.
  1. (1)
    If $N$ has exactly one cusp, then $N$ is either homeomorphic to $\mathcal{L}(n-\frac{1}{2})$ for some $n\in\mathbb{Z}$, or homeomorphic to $L_{B}(0,r)$ for some $r\in\mathbb{Q}\setminus\{0,4\}$.
  2. (2)
    If $N$ has more than one cusp, then $N$ is homeomorphic to $L_{B}(0)$.
Since $\infty,-2,-1,0,1,2$ are all the exceptional slopes on one component of $\mathcal{L}$ according to [37], $\mathcal{L}(n-\frac{1}{2})$ is hyperbolic for all $n\in\mathbb{Z}$. Furthermore, [32, Proposition 2.6] showed that $L_{B}(0)$ is hyperbolic, and [32, Proposition 2.7] showed that $L_{B}(0,r)$ is hyperbolic for any $r\in\mathbb{Q}\setminus\{0,4\}$.
In the following lemma, we list in details the boundary slopes of the manifolds in Proposition 6.17 yielding Klein bottles that achieves the geometric intersection number $4$, and specify the corresponding Dehn filled 3-manifolds.
Recall that $\mathcal{N}$ is the orientable $I$-bundle over Klein bottle, and $M_{1}$ is the complement of the Whitehead link $\mathcal{W}$ in $S^{3}$. In the following context, we denote $M_{3_{1}}$ as the complement of the trefoil in $S^{3}$, i.e. $M_{3_{1}}$ is a Seifert fibered space $(D^{2};(2,1),(3,1))$.
Lemma 6.18.
  1. (1)
    $L_{B}(0,0)\cong L_{B}(0,4)$ can be decomposed as $\mathcal{N}\cup_{T^{2}}M_{1}$, i.e. gluing $\mathcal{N}$ and $M_{1}$ along a torus boundary.
  2. (2)
    For each $r\in\mathbb{Q}$, $L_{B}(0,r,0)$ and $L_{B}(0,r,4)$ contain embedded Klein bottles. They can be decomposed as $\mathcal{N}\cup_{T^{2}}M^{\prime}$, where $\partial M^{\prime}=T^{2}$ and $M^{\prime}$ is either a hyperbolic manifold, a Seifert fibered space with hyperbolic base orbifold, or a graph manifold with non-trivial JSJ-decomposition.
  3. (3)
    $\mathcal{L}(\frac{3}{2},-2)=(\mathbb{RP}^{2};(2,1),(3,1),-1)$ and $\mathcal{L}(n-\frac{1}{2},-2)=\mathcal{N}\cup_{T^{2}}M_{3_{1}}$ for $n\in\mathbb{Z}\setminus\{2\}$.
  4. (4)
    $\mathcal{L}(-\frac{3}{2},2)=(\mathbb{RP}^{2};(2,1),(3,1),-1)$ and $\mathcal{L}(n-\frac{1}{2},2)=\mathcal{N}\cup_{T^{2}}M_{3_{1}}$ for $n\in\mathbb{Z}\setminus\{-1\}$.
In particular, all the above listed Dehn-filled manifolds are aspherical, and contain an embedded Klein bottle.
Proof.
(1) is proved in [32, Lemma 2.3 (2)].
According to (6.15), $L_{B}(0,r,0)\cong L_{B}(0,0,r)$ and $L_{B}(0,r,4)\cong L_{B}(0,4,r)$. Thus, they can be viewed as Dehn fillings on $L_{B}(0,0)$ and $L_{B}(0,4)$. According to (1), $L_{B}(0,0,r)$ and $L_{B}(0,4,r)$ can be decomposed by an embedded torus into $\mathcal{N}\cup M^{\prime}$, where $M^{\prime}$ is a Dehn filling of $M_{1}$ on one of its boundary components $\partial_{1}M_{1}$. The Dehn fillings of $M_{1}$ on $\partial_{1}M_{1}$ have been classified by [37, Table A.1]. In fact, there is exactly one slope yielding the solid torus $D^{2}\times S^{1}$, while the other slopes yield either a hyperbolic manifold, a Seifert fibered space with hyperbolic base orbifold, or a graph manifold with non-trivial JSJ-decomposition. In other words, there is only one element $r_{0}\in\mathbb{Q}\cup\{\infty\}$ such that $L_{B}(0,0,r_{0})\cong\mathcal{N}\cup_{T^{2}}(D^{2}\times S^{1})$, which is indeed non-aspherical; while for any $r\neq r_{0}\in\mathbb{Q}\cup\{\infty\}$, $L_{B}(0,0,r)\cong\mathcal{N}\cup_{T^{2}}M^{\prime}$ is aspherical. We claim that $r_{0}=\infty$. In fact, by (6.15), $L_{B}(0,0,\infty)\cong L_{B}(0,\infty,0)$ is a Dehn filling on $L_{B}(0,\infty)$. [32, Lemma 2.3 (1)] showed that $L_{B}(0,\infty)\cong\mathcal{N}$, and thus $L_{B}(0,\infty,0)\cong\mathcal{N}\cup_{T^{2}}(D^{2}\times S^{1})$. Similarly, there is only one element $r_{4}\in\mathbb{Q}\cup\{\infty\}$ such that $L_{B}(0,4,r_{4})\cong\mathcal{N}\cup_{T^{2}}(D^{2}\times S^{1})$ is spherical, while for any $r\neq r_{4}\in\mathbb{Q}\cup\{\infty\}$, $L_{B}(0,4,r_{4})\cong\mathcal{N}\cup_{T^{2}}M^{\prime}$ is aspherical. One can similarly show that $r_{4}=\infty$ by [32, Lemma 2.3 (1)]. Thus, this proves (2).
(3) follows from [37, Table 3]. In fact, the link $\mathcal{L}$ is obtained by a $-\frac{1}{2}$-surgery on one component of the three-chain link Figure 10 in $S^{3}$. Thus, for any $r,s\in\mathbb{Q}\cup\{\infty\}$, $\mathcal{L}(r,s)$ can be viewed as $(-\frac{1}{2},r-2,s-2)$-surgery on three components of the three-chain link, for which [37] presents a complete classification of exceptional Dehn surgeries.
Figure 10. The three-chain
(4) follows directly from (3) together with the achirality of $\mathcal{L}$ (6.16).
Note that the above listed Dehn-filled manifolds are either irreducible 3-manifolds admitting non-trivial JSJ-decomposition, or Seifert fibered spaces with hyperbolic base orbifolds. Thus, these Dehn-filled manifolds are all aspherical. They obviously contain embedded Klein bottles, as they satisfy Lemma 5.6 (ii). ∎
Recall that in the definition of the subset $\mathcal{E}_{k}(M,\partial_{i_{0}}M)$ of exceptional slopes, we require the Dehn-filled manifold which contains Klein bottles to be aspherical. Therefore, by Lemma 6.18, we can obtain the following corollary.
Corollary 6.19.
We follow the ${}\mathbb{Q}\cup\{\infty\}$-parametrization of boundary slopes of the link complements $S^{3}\setminus\mathring{N}(\mathcal{L})$ and $S^{3}\setminus\mathring{N}(L_{B})$ as the parametrization for the boundary slopes of $L_{B}(0)$, $L_{B}(0,r)$ and $\mathcal{L}(n-\frac{1}{2})$. Then,
  1. (1)
    $\{0,4\}\subseteq\mathcal{E}_{k}\left(L_{B}(0),\partial N(L_{B}^{(2)})\right)$, and $\Delta(0,4)=4$;
  2. (2)
    $\{0,4\}\subseteq\mathcal{E}_{k}\left(L_{B}(0,r),\partial N(L_{B}^{(3)})\right)$ for each $r\in\mathbb{Q}\setminus\{0,4\}$, and $\Delta(0,4)=4$;
  3. (3)
    $\{-2,2\}\subseteq\mathcal{E}_{k}\left(\mathcal{L}(n-\frac{1}{2}),\partial(\mathcal{L}(n-\frac{1}{2}))\right)$ for each $n\in\mathbb{Z}$, and $\Delta(-2,2)=4$.
With the asphericity of the Dehn-filled manifolds verified, we can deduce profinite rigidity of these manifolds from Proposition 5.10 (4).
Theorem 6.20.
$L_{B}(0)$ is profinitely rigid among all compact, orientable 3-manifolds.
Proof.
Suppose $N$ is a compact orientable 3-manifold such that $\widehat{\pi_{1}L_{B}(0)}\cong\widehat{\pi_{1}N}$. Recall that $L_{B}(0)$ is finite-volume hyperbolic by [32, Proposition 2.6]. Thus, $N$ is also finite-volume hyperbolic with two cusps according to Proposition 2.2.
By Corollary 6.19 (1), there exist two slopes $c_{1},c_{2}\in\mathcal{E}_{k}(L_{B}(0),\partial(L_{B}(0)))$ such that $\Delta(c_{1},c_{2})=4$. Then, Proposition 5.10 together with Lemma 6.2 implies that there exist two slopes $c_{1}^{\prime},c_{2}^{\prime}\in\mathcal{E}_{k}(N,\partial_{j}N)$ such that $\Delta(c_{1}^{\prime},c_{2}^{\prime})=4$. Since $N$ has two cusps, Proposition 6.17 implies that $N$ is homeomorphic to $L_{B}(0)$. ∎
To prove the profinite rigidity of $\mathcal{L}(n-\frac{1}{2})$ and $L_{B}(0,r)$, we have to firstly distinguish these manifolds through the profinite completion of their fundamental groups. Let us first make some preparations.
Lemma 6.21.
Let $M$ be one of $\mathcal{L}(n-\frac{1}{2})$ ($n\in\mathbb{Z}$) or $L_{B}(0,r)$ ($r\in\mathbb{Q}\setminus\{0,4\}$). Then $\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{k}(M,\partial M)\}=4$.
Proof.
We have shown by Corollary 6.19 that $\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{k}(M,\partial M)\}\geq 4$. If $\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{k}(M,\partial M)\}>4$, then [33, Corollary 1.5] shows that $M$ is homeomorphic to one of $\mathcal{W}(5),\mathcal{W}(-1),\mathcal{W}(-2),\mathcal{W}(-4)$. Thus, it suffices to show that $M$ is not homeomorphic to $\mathcal{W}(5),\mathcal{W}(-1),\mathcal{W}(-2),\mathcal{W}(-4)$.
In fact, according to [37, Table A.2] for $\mathcal{W}(5),\mathcal{W}(-1),\mathcal{W}(-2)$ and [37, Table A.4] for $\mathcal{W}(-4)$:
  1. (i)
    $\mathcal{E}_{k}(\mathcal{W}(5),\partial\mathcal{W}(5))=\{c_{5},c_{5}^{\prime}\}$, and $\Delta(c_{5},c_{5}^{\prime})=8$;
  2. (ii)
    $\mathcal{E}_{k}(\mathcal{W}(-1),\partial\mathcal{W}(-1))=\{c_{-1},c_{-1}^{\prime}\}$, and $\Delta(c_{-1},c_{-1}^{\prime})=8$;
  3. (iii)
    $\mathcal{E}_{k}(\mathcal{W}(-2),\partial\mathcal{W}(-2))=\{c_{-2},c_{-2}^{\prime}\}$, and $\Delta(c_{-2},c_{-2}^{\prime})=6$;
  4. (iv)
    $\mathcal{E}_{k}(\mathcal{W}(-4),\partial\mathcal{W}(-4))=\{c_{-4},c_{-4}^{\prime}\}$, and $\Delta(c_{-4},c_{-4}^{\prime})=5$.
None of them contain two slopes with geometric intersection number $4$. Therefore, by Corollary 6.19, $M$ is not homeomorphic to $\mathcal{W}(5),\mathcal{W}(-1),\mathcal{W}(-2),\mathcal{W}(-4)$, finishing the proof. ∎
Lemma 6.22.
Let $T$ and $T^{\prime}$ be two tori.
  1. (1)
    Suppose $E\subseteq\text{slp}(T)$ is a subset such that $\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in E\}=4$. Then the pair $c_{1},c_{2}\in E$ realizing the maximal geometric intersection number $\Delta(c_{1},c_{2})=4$ is unique.
  2. (2)
    Let $c_{1},c_{2}\in\text{slp}(T)$ and $c_{1}^{\prime},c_{2}^{\prime}\in\text{slp}(T^{\prime})$ such that $\Delta(c_{1},c_{2})=\Delta(c_{1}^{\prime},c_{2}^{\prime})=4$.
    Then there exist exactly two (equivalence classes of) isomorphisms $\psi_{1},\psi_{2}\in\mathrm{Isom}_{\mathbb{Z}}(\pi_{1}(T),\pi_{1}(T^{\prime}))/(\psi\sim-\psi)\cong\mathrm{PGL}(2,\mathbb{Z})$ such that $\overline{\psi}_{i}(\{c_{1},c_{2}\})=\{c_{1}^{\prime},c_{2}^{\prime}\}$.
Proof.
(1): Suppose $c_{1},c_{2}\in E$ such that $\Delta(c_{1},c_{2})=4$. We can choose a basis $(\mu_{0},\lambda_{0})$ for $\pi_{1}(T)\cong\mathbb{Z}^{2}$, so that $c_{1}=[\mu_{0}]$. Then, $c_{2}=[a\mu_{0}+4\lambda_{0}]$ for some odd integer $a$. We can assume $a=4k+t$, where $k\in\mathbb{Z}$ and $t\in\{\pm 1\}$. Let $\mu=t\mu_{0}$ and $\lambda=k\mu_{0}+\lambda_{0}$. Then $(\mu,\lambda)$ is also a basis for $\pi_{1}(T)$, and $c_{1}=[\mu]$, $c_{2}=[\mu+4\lambda]$.
We can now list all possible elements in $E$. Suppose $c=[p\mu+q\lambda]\in E$, where $(p,q)$ is a pair of coprime integers. Without loss of generality, we may assume $p\geq 0$. Then $\Delta(c_{1},c)=\left|q\right|\leq 4$, and $\Delta(c_{2},c)=\left|4p-q\right|\leq 4$. It is easy to verify that $(p,q)\in\{(0,\pm 1),(1,0),(1,1)(1,2),(1,3),(1,4)\}$. Thus, $E\subseteq\{[\mu],[\mu+\lambda],[\mu+2\lambda],[\mu+3\lambda],[\mu+4\lambda],[\lambda]\}$, and it is easy to verify that $[\mu]$ and $[\mu+4\lambda]$ is the unique pair realizing geometric intersection number $4$.
(2): According to (1), we can choose basis $(\mu,\lambda)$ for $\pi_{1}(T)$, and $(\mu^{\prime},\lambda^{\prime})$ for $\pi_{1}(T^{\prime})$, such that $c_{1}=[\mu]$, $c_{2}=[\mu+4\lambda]$, $c_{1}^{\prime}=[\mu^{\prime}]$, and $c_{2}^{\prime}=[\mu^{\prime}+4\lambda^{\prime}]$. Suppose $\psi:\pi_{1}(T)\to\pi_{1}(T^{\prime})$ is an isomorphism such that $\overline{\psi}(\{c_{1},c_{2}\})=\{c_{1}^{\prime},c_{2}^{\prime}\}$. There are two cases to be considered.
In the first case, $\overline{\psi}(c_{1})=c_{1}^{\prime}$ and $\overline{\psi}(c_{2})=c_{2}^{\prime}$. The only possibilities are $\psi(\mu,\mu+4\lambda)=\pm(\mu^{\prime},\mu^{\prime}+4\lambda^{\prime})$, since otherwise $4\psi(\lambda)=\pm(2\mu^{\prime}+4\lambda^{\prime})$, which is impossible. Thus, $\psi(\mu,\lambda)=\pm(\mu^{\prime},\lambda^{\prime})$.
In the second case, $\overline{\psi}(c_{1})=c_{2}^{\prime}$ and $\overline{\psi}(c_{2})=c_{1}^{\prime}$. The only possibilities are $\psi(\mu,\mu+4\lambda)=\pm(\mu^{\prime}+4\lambda^{\prime},\mu^{\prime})$, since otherwise we again obtain $4\psi(\lambda)=\pm(2\mu^{\prime}+4\lambda^{\prime})$, which is impossible. Thus, $\psi(\mu,\lambda)=\pm(\mu^{\prime}+4\lambda^{\prime},-\lambda^{\prime})$.
Therefore, there are altogether two isomorphisms $\psi_{1},\psi_{2}\in\mathrm{Isom}_{\mathbb{Z}}(\pi_{1}(T),\pi_{1}(T^{\prime}))/(\psi\sim-\psi)$ such that $\overline{\psi}_{i}(\{c_{1},c_{2}\})=\{c_{1}^{\prime},c_{2}^{\prime}\}$. ∎
Lemma 6.23.
Suppose $M$ and $N$ belong to $\mathcal{L}(n-\frac{1}{2})$ ($n\in\mathbb{Z}$) or $L_{B}(0,r)$ ($r\in\mathbb{Q}\setminus\{0,4\}$). Then $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$ if and only if $M\cong N$.
Proof.
We first prove that $\mathcal{L}(n-\frac{1}{2})$ is not profinitely isomorphic to $L_{B}(0,r)$, by showing that their first homologies are not isomorphic. In fact, up to an appropriate choice of orientations, the two components of $\mathcal{L}$ have linking number $3$. Let $\alpha,\beta$ be two generators of $H_{1}(S^{3}\setminus\mathring{N}(\mathcal{L});\mathbb{Z})\cong\mathbb{Z}^{2}$ corresponding to the meridians of the two components. Then,
$H_{1}(\mathcal{L}(n-\frac{1}{2});\mathbb{Z})\cong\frac{H_{1}(S^{3}\setminus\mathring{N}(\mathcal{L});\mathbb{Z})}{\left\langle(2n-1)\alpha+6\beta\right\rangle}\cong\left\{\begin{aligned} &\mathbb{Z}\oplus\mathbb{Z}/3\mathbb{Z},&3\mid 2n-1\\ &\mathbb{Z},&3\nmid 2n-1\end{aligned}\right..$
Similarly, up to an appropriate choice of orientations, the three components $L_{B}^{(1)},L_{B}^{(2)},L_{B}^{(3)}$ have pairwise linking number $2$. Let $\alpha,\beta,\gamma$ be three generators of $H_{1}(S^{3}\setminus\mathring{N}(L_{B}))\cong\mathbb{Z}^{3}$ corresponding to the meridians of $L_{B}^{(1)},L_{B}^{(2)},L_{B}^{(3)}$ respectively. For $r=\frac{p}{q}\in\mathbb{Q}$ a reduced fraction,
(6.24) $H_{1}(L_{B}(0,r);\mathbb{Z})\cong\frac{H_{1}(S^{3}\setminus\mathring{N}(L_{B});\mathbb{Z})}{\left\langle 2\beta+2\gamma,2q\alpha+p\beta+2q\gamma\right\rangle}\cong\left\{\begin{aligned} &\mathbb{Z}\oplus\mathbb{Z}/2\mathbb{Z}\oplus\mathbb{Z}/2\mathbb{Z},&2\mid p\\ &\mathbb{Z}\oplus\mathbb{Z}/2\mathbb{Z},&2\nmid p\end{aligned}\right..$
Therefore, $H_{1}(\mathcal{L}(n-\frac{1}{2});\mathbb{Z})\not\cong H_{1}(L_{B}(0,r);\mathbb{Z})$, which implies $\widehat{\pi_{1}\mathcal{L}(n-\frac{1}{2})}\not\cong\widehat{\pi_{1}L_{B}(0,r)}$.
Now we consider the family $\mathcal{L}(n-\frac{1}{2})$ ($n\in\mathbb{Z}$). Let $M=\mathcal{L}(m-\frac{1}{2})$ and $N=\mathcal{L}(n-\frac{1}{2})$, and suppose $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$. Then according to Theorem A, there exists an isomorphism $\psi:\pi_{1}\partial M\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}\partial N$ such that $\widehat{\pi_{1}M_{c}}\cong\widehat{\pi_{1}N_{\overline{\psi}(c)}}$ for all $c\in\text{slp}(\partial M)$. In particular, $\overline{\psi}(\mathcal{E}_{k}(M,\partial M))=\mathcal{E}_{k}(N,\partial N)$ according to Proposition 5.10. We follow the ${}\mathbb{Q}\cup\{\infty\}$-parametrization for $\text{slp}(\partial M)$ and $\text{slp}(\partial N)$ as above. Recall that $\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{k}(M,\partial M)\}=\max\{\Delta(c_{1},c_{2})\mid c_{1},c_{2}\in\mathcal{E}_{k}(N,\partial N)\}=4$ by Lemma 6.21. Thus, according to Corollary 6.19 (3) and Lemma 6.22 (1), $\{-2,2\}$ is the unique pair of elements in $\mathcal{E}_{k}(M,\partial M)$, as well as in $\mathcal{E}_{k}(N,\partial N)$, that realizes the maximal geometric intersection number $4$. Since $\overline{\psi}:\mathcal{E}_{k}(M,\partial M)\xrightarrow{1:1}\mathcal{E}_{k}(N,\partial N)$ preserves the geometric intersection number by Lemma 6.2, we derive that $\overline{\psi}(\{-2,2\})=\{-2,2\}$. According to Lemma 6.22 (2), there are only two possible candidates for $\psi$ in $\mathrm{Isom}_{\mathbb{Z}}(\pi_{1}\partial M,\pi_{1}\partial N)/(\psi\sim-\psi)$. Indeed, under the ${}\mathbb{Q}\cup\{\infty\}$-parametrization, either $\overline{\psi}(s)=s$, or $\overline{\psi}(s)=-s$ for all $s\in\mathbb{Q}\cup\{\infty\}$.
In addition, $\overline{\psi}$ sends the rational null-homologous boundary slope of $M$ to the rational null-homologous boundary slope of $N$ by Lemma 4.4 (2). Since the linking number of the two components of $\mathcal{L}$ is $3$, the rational null-homologous boundary slope of $M$ corresponds to $\frac{3^{2}}{m-\frac{1}{2}}=\frac{18}{2m-1}$ in the ${}\mathbb{Q}\cup\{\infty\}$-parametrization. Similarly, the rational null-homologous boundary slope of $N$ corresponds to $\frac{3^{2}}{m-\frac{1}{2}}=\frac{18}{2n-1}$. Thus, $\overline{\psi}(\frac{18}{2m-1})=\frac{18}{2n-1}$. Recall that $\overline{\psi}$ is a multiplication by ±$\pm 1$ under the ${}\mathbb{Q}\cup\{\infty\}$-parametrization. Thus, $2m-1=\pm(2n-1)$. Note that $\mathcal{L}(r)\cong\mathcal{L}(-r)$ according to (6.16). Thus, regardless of the $\pm$-sign, $M=\mathcal{L}(\frac{2m-1}{2})\cong N=\mathcal{L}(\frac{2n-1}{2})$.
Next, we consider the family $L_{B}(0,r)$ ($r\in\mathbb{Q}\setminus\{0,4\}$). Let $M=L_{B}(0,r)$ and $N=L_{B}(0,r^{\prime})$ and suppose $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$. Similarly, by Theorem A, there exists an isomorphism $\psi:\pi_{1}\partial M\to\pi_{1}\partial N$ such that $\widehat{\pi_{1}M_{c}}\cong\widehat{\pi_{1}N_{\overline{\psi}(c)}}$, and $\overline{\psi}(\mathcal{E}_{k}(M,\partial M))=\mathcal{E}_{k}(N,\partial N)$ according to Proposition 5.10. Again by Lemma 6.21, Corollary 6.19 (2) and Lemma 6.22 (1), $\{0,4\}$ is the only pair of elements in $\mathcal{E}_{k}(M,\partial M)$, as well as in $\mathcal{E}_{k}(N,\partial N)$, that realizes the maximal geometric intersection number $4$, and Lemma 6.2 implies that $\overline{\psi}(\{0,4\})=\{0,4\}$. According to Lemma 6.22 (2), there are only two possible candidates for $\psi$ up to $\pm$-sign; under the ${}\mathbb{Q}\cup\{\infty\}$-parametrization, either $\overline{\psi}(u)=u$, or $\overline{\psi}(u)=4-u$ for all $u\in\mathbb{Q}\cup\{\infty\}$.
We still consider the rational null-homologous boundary slopes. Let $r=\frac{p}{q}$ be a reduced fraction. Then (6.24) implies that $2q\alpha+2q\beta+(4q-p)\gamma$ is torsion in $H_{1}(L_{B}(0,r);\mathbb{Z})$. Thus, the slope represented by $\frac{4q-p}{q}=4-r$ on $\partial N(L_{B}^{(3)})$ is the rational null-homologous boundary slope on $\partial M$. Similarly, $4-s$ is the rational null-homologous boundary slope on $\partial N$. By Lemma 4.4 (2), $\overline{\psi}(4-r)=4-s$. Recall that either $\overline{\psi}(u)=u$, or $\overline{\psi}(u)=4-u$. Thus, either $s=r$ or $s=4-r$. Fortunately, [32, Lemma 2.2 (2)] proved that $L_{B}(0,r)\cong L_{B}(0,4-r)$. Thus, in both cases, $M\cong N$. ∎
Finally, we are ready to prove the profinite rigidity of these two families.
Theorem 6.25.
Let $M$ be one of $\mathcal{L}(n-\frac{1}{2})$ ($n\in\mathbb{Z}$) or $L_{B}(0,r)$ ($r\in\mathbb{Q}\setminus\{0,4\}$). Then $M$ is profinitely rigid among all compact, orientable 3-manifolds.
Proof.
Suppose $N$ is a compact, orientable 3-manifold such that $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$. Recall that $M$ is hyperbolic by [37] and [32, Proposition 2.7]. Then according to Proposition 2.2, $N$ is also finite-volume hyperbolic with exactly one cusp.
By Corollary 6.19, there exist two slopes $c_{1},c_{2}\in\mathcal{E}_{k}(M,\partial M)$ such that $\Delta(c_{1},c_{2})=4$. Then, Proposition 5.10 together with Lemma 6.2 implies that there exist two slopes $c_{1}^{\prime},c_{2}^{\prime}\in\mathcal{E}_{k}(N,\partial_{j}N)$ such that $\Delta(c_{1}^{\prime},c_{2}^{\prime})=4$. Since $N$ has exactly one cusp, Proposition 6.17 implies that $N$ is either homeomorphic to $\mathcal{L}(n-\frac{1}{2})$ for some $n\in\mathbb{Z}$, or to $L_{B}(0,r)$ for some $r\in\mathbb{Q}\setminus\{0,4\}$. It then follows from Lemma 6.23 that $N\cong M$, finishing the proof. ∎
Appendix A Proof of Proposition 3.3
Lemma A.1.
Let $M$ be a compact, orientable, irreducible and boundary-incompressible 3-manifold. Suppose $N$ is a compact, orientable 3-manifold such that $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$. Then,
  1. (1)
    $N$ is also irreducible and boundary-incompressible;
  2. (2)
    if $M$ is closed, then $N$ is also closed;
  3. (3)
    if $\partial M$ consists of tori, then so does $\partial N$.
Proof.
As a standard application of the loop theorem, through attaching 2-handles to the compressible boundary components of $N$ and then capping off the emerging boundary spheres by 3-handles, we can obtain a boundary-incompressible 3-manifold $N^{+}$ such that $\pi_{1}N\cong\pi_{1}N^{+}$. Then $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N^{+}}$.
[51, Theorem 6.22] showed that the profinite completion of fundamental group determines the prime decomposition of a compact, orientable, boundary-incompressible 3-manifold. In particular, when $M$ is irreducible, $N^{+}$ is also irreducible.
Note that when $N$ is boundary-compressible, attaching a 2-handle along a compressible boundary curve will create an essential sphere, which makes $N^{+}$ reducible. Thus, $N$ is boundary-incompressible and $N=N^{+}$. Therefore, $N$ is also irreducible, which proves (1). (2) and (3) then follows from [57, Corollary 4.1]. ∎
Before moving on, let us briefly recall some useful concepts related to JSJ-decomposition.
Let $M$ be a compact, orientable, irreducible 3-manifold with empty or incompressible toroidal boundary, which is not a closed $Sol$-manifold. Then there exists a minimal collection $\mathcal{T}$ of disjoint essential tori, such that each connected component of $M\setminus\mathcal{T}$ is either a Seifert fibered manifold or a finite-volume hyperbolic manifold. In addition, $\mathcal{T}$ is unique up to isotopy, and is called the JSJ-tori of $M$.
The JSJ-decomposition endows $M$ with a structure of a graph of spaces, so that it splits $\pi_{1}M$ as the fundamental group of a graph of groups: $\pi_{1}M=\pi_{1}(\mathcal{G},\Gamma_{M})$. Here, $\Gamma_{M}$ denotes the JSJ-graph of $M$, consisting of vertices $V(\Gamma_{M})$ and edges $E(\Gamma_{M})$; for each $v\in V(\Gamma_{M})$, $\mathcal{G}_{v}$ is the fundamental group of a JSJ-piece (either hyperbolic or Seifert), and for each $e\in E(\Gamma_{M})$, $\mathcal{G}_{e}=\pi_{1}(T^{2})\cong\mathbb{Z}\oplus\mathbb{Z}$. Figure 11 shows an example for this construction and we refer the readers to [43] for details.
(a) The irreducible 3-manifold $M$
$u$$v$$w$
(b) The JSJ-graph $\Gamma_{M}$
$\pi_{1}M_{u}$$\pi_{1}T^{2}$$\pi_{1}M_{v}$$\pi_{1}T^{2}$$\pi_{1}M_{w}$$\pi_{1}T^{2}$
(c) The graph of groups $(\mathcal{G}_{M},\Gamma_{M})$
Figure 11. JSJ-graph of groups
Proposition A.2 ([53, Theorem A] ).
Let $M$ be compact, orientable, irreducible 3-manifold with empty or incompressible toral boundary. For each $x\in\Gamma_{M}=V(\Gamma_{M})\cup E(\Gamma_{M})$, $\pi_{1}M$ induces the full profinite topology on $\mathcal{G}_{x}$. Thus, the homomorphism $\widehat{\mathcal{G}_{x}}\to\widehat{\pi_{1}M}$ is injective.
The proof of [53, Theorem A] was stated for closed manifolds, while it applies to manifolds with boundary with no alteration.
Proposition A.3 ([56, Theorem 4.3] ).
Let $M$ and $N$ be compact, orientable, irreducible 3-manifolds with empty or incompressible toroidal boundary, which are not closed $Sol$-manifolds. Let $(\mathcal{G},\Gamma_{M})$ and $(\mathcal{H},\Gamma_{N})$ denote their JSJ-graphs of groups respectively. Suppose $f:\widehat{\pi_{1}M}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N}$ is an isomorphism.
  1. (1)
    There is a graph isomorphism $f^{\dagger}:\Gamma_{M}\to\Gamma_{N}$, which preserves the hyperbolic/Seifert type of vertices.
  2. (2)
    For each $x\in\Gamma_{M}$, $f$ induces an isomorphism $f_{x}:\widehat{\mathcal{G}_{x}}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\mathcal{H}_{f^{\dagger}(x)}}$.
  3. (3)
    For each $x\in\Gamma_{M}$, there is an element $h_{x}\in\widehat{\pi_{1}N}$ such that the following diagram commutes, where $C_{h_{x}}$ denotes the conjugation by $h_{x}$.
    ${\widehat{\mathcal{G}_{x}}}$${\widehat{\mathcal{H}_{f^{\dagger}(x)}}}$${\widehat{\pi_{1}M}}$${\widehat{\pi_{1}N}}$${\widehat{\pi_{1}N}}$$\scriptstyle{f_{x}}$$\scriptstyle{f}$$\scriptstyle{C_{h_{x}}}$
In fact, [56] was originally proven for closed manifolds, but the proof can be easily generalized to the bounded case, see [48, Theorem 5.3] or [57, Theorem 5.6].
We are now ready to prove Proposition 3.3.
Proposition 3.3.
Let $M$ be a mixed 3-manifold, and let $\partial_{1}^{hyp}M,\cdots,\partial_{n}^{hyp}M$ be the boundary components of $M$ belonging to a hyperbolic JSJ-piece. Suppose $N$ is a compact, orientable 3-manifold, and $f:\widehat{\pi_{1}M}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N}$ is an isomorphism.
  1. (1)
    $N$ is also a mixed 3-manifold.
  2. (2)
    There are exactly $n$ boundary components of $N$ which belong to hyperbolic JSJ-pieces, and we denote these boundary components as $\partial_{1}^{hyp}N,\cdots,\partial_{n}^{hyp}N$.
  3. (3)
    Up to a reordering, $f(\overline{\pi_{1}\partial_{i}^{hyp}M})$ is a conjugate of $\overline{\pi_{1}\partial_{i}^{hyp}N}$ in $\widehat{\pi_{1}N}$.
  4. (4)
    $f$ is peripheral $\widehat{\mathbb{Z}}^{\times}$-regular at each $\partial_{i}^{hyp}M$ ($1\leq i\leq n$).
Proof.
(1) By Lemma A.1, $N$ is irreducible and has empty or incompressible toral boundary. In addition, by [54, Theorem 8.4], $N$ is not a closed $Sol$-manifold. Therefore, $N$ satisfy the requirements for Proposition A.3. In particular, Proposition A.3 (1) implies that $N$ is also a mixed 3-manifold.
(2) We follow the notation in Proposition A.3. $(\mathcal{G},\Gamma_{M})$ and $(\mathcal{H},\Gamma_{N})$ denote the JSJ-graphs of groups, $f^{\dagger}:\Gamma_{M}\to\Gamma_{N}$ denotes the graph isomorphism, and $f_{x}:\widehat{\mathcal{G}_{x}}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\mathcal{H}_{f^{\dagger}(x)}}$ denotes the induced isomorphism for each $x\in\Gamma_{M}$. For each vertex $v\in V(\Gamma_{M})$ corresponding to a hyperbolic piece, according to Proposition 2.2, $\widehat{\pi_{1}M_{v}}\cong\widehat{\pi_{1}N_{f^{\dagger}(v)}}$ implies that $M_{v}$ and $N_{f^{\dagger}(v)}$ have the same number of boundary components. Let $\partial^{hyp}M$ and $\partial^{hyp}N$ denote the boundary components belonging to hyperbolic pieces. Together with the graph isomorphism $f^{\dagger}:\Gamma_{M}\to\Gamma_{N}$,
$n=\sharp\partial^{hyp}M=\sum_{{v\in V(\Gamma_{M})}\atop{\text{hyperbolic}}}\left(\sharp\partial M_{v}-deg_{\Gamma_{M}}(v)\right)=\sum_{{w\in V(\Gamma_{N})}\atop{\text{hyperbolic}}}\left(\sharp\partial N_{w}-deg_{\Gamma_{N}}(w)\right)=\sharp\partial^{hyp}N.$
(3) For each hyperbolic vertex $v\in V(\Gamma_{M})$, $f_{v}:\widehat{\pi_{1}M_{v}}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N_{f^{\dagger}(v)}}$ respects the peripheral structure according to [57, Lemma 6.3] (or Theorem 3.2 (1)). In fact, according to the commutative diagram in Proposition A.3 (3), $f_{v}$ matches up the boundary components of $M_{v}$ coming from the JSJ-tori (i.e. corresponding to the edges $e\in E(\Gamma_{M})$ adjoining $v$) to the boundary components of $N_{f^{\dagger}(v)}$ coming from the JSJ-tori (i.e. corresponding to the edges $f^{\dagger}(e)$). Thus, $f_{v}$ matches up the remaining boundary components, i.e. matches up $M_{v}\cap\partial M$ with $N_{f^{\dagger}(v)}\cap\partial N$. Therefore, up to reordering, $f_{v}$ sends $\overline{\pi_{1}\partial_{i}^{hyp}M}$ to a conjugate of $\overline{\pi_{1}\partial_{i}^{hyp}N}$ in $\widehat{\pi_{1}N_{f^{\dagger}(v)}}$. The commutative diagram in Proposition A.3 (3) then implies that $f$ in fact sends $\overline{\pi_{1}\partial_{i}^{hyp}M}$ to a conjugate of $\overline{\pi_{1}\partial_{i}^{hyp}N}$ in $\widehat{\pi_{1}N}$.
(4) Based on our proof for (3), (4) follows from the fact that $f_{v}$ is peripheral $\widehat{\mathbb{Z}}^{\times}$-regular at $\partial_{i}^{hyp}M$ (Theorem 3.2 (2)), together with the commutative diagram Proposition A.3 (3). ∎